### v1

# Orient Anything

Abstract
Orientation estimation is a fundamental task in 3D shape analysis which consists of estimating a shape’s orientation axes: its side-, up-, and front-axes. Using this data, one can rotate a shape into canonical orientation, where its orientation axes are aligned with the coordinate axes. Developing an orientation algorithm that reliably estimates complete orientations of general shapes remains an open problem. We introduce a two-stage orientation pipeline that achieves state of the art performance on up-axis estimation and further demonstrate its efficacy on full-orientation estimation, where one seeks all three orientation axes. Unlike previous work, we train and evaluate our method on all of Shapenet rather than a subset of classes. We motivate our engineering contributions by theory describing fundamental obstacles to orientation estimation for rotationally-symmetric shapes, and show how our method avoids these obstacles.
1 Introduction
Orientation estimation is a fundamental task in 3D shape analysis which consists of estimating a shape’s orientation axes: its side-, up-, and front-axes. Using this data, one can rotate a shape into canonical orientation, in which the shape’s orientation axes are aligned with the coordinate axes. This task is especially important as a pre-processing step in 3D deep learning, where deep networks are typically trained on datasets of canonically-oriented shapes but applied to arbitrarily-oriented shapes at inference time. While data augmentation or equivariant and invariant architectures may improve a model’s robustness to input rotations, these techniques come at the cost of data efficiency and model expressivity (Kuchnik & Smith, 2019; Kim et al., 2023). In contrast, orientation estimation allows one to pre-process shapes at inference time so that their orientation matches a model’s training data.
Orientation estimation is a challenging task, and developing an orientation pipeline that reliably estimates complete orientations of general shapes remains an open problem. The naïve deep learning approach is to train a model with an $L^{2}$ loss to directly predict a shape’s orientation from a point cloud of surface samples. However, this strategy fails for shapes with rotational symmetries, where the optimal solution to the $L^{2}$ regression problem is the Euclidean mean (Moakher, 2002) of a shape’s orientations over all of its symmetries. In contrast, works such as Poursaeed et al. (2020) discretize the unit sphere into a set of fixed rotations and train a classifier to predict a probability distribution over these rotations, but find that this approach fails for any sufficiently dense discretization of the unit sphere.
Our key insight is to divide orientation estimation into two tractable sub-problems. In the first stage (the quotient orienter), we solve a continuous regression problem to recover a shape’s orientation up to octahedral symmetries. In the second stage (the flipper), we solve a discrete classification problem to predict one of 24 octahedral flips that returns the first-stage output to canonical orientation. Octahedral symmetries form a small set covering a substantial proportion of the symmetries occurring in real-world shapes. Consequently, quotienting our first-stage regression problem by octahedral symmetries prevents its predictions from collapsing to averages, while also keeping the subsequent classification problem tractable.
Using this strategy, our method achieves state-of-the-art performance on the well-studied problem of up-axis prediction, and additionally performs well on full-orientation prediction, which few prior works have tackled. Unlike previous work, we train and evaluate our model on the entire Shapenet dataset rather than a subset of classes. We further demonstrate its generalization capabilities on Objaverse, a large dataset of real-world 3D models of varying quality.
A shape’s ground truth orientation may be ambiguous. This challenge is especially salient for nearly-symmetric shapes, where multiple orientations may yield nearly indistinguishable shapes. To resolve this issue, we use conformal prediction to enable our flipper to output adaptive prediction sets (Romano et al., 2020) whose size varies with the flipper model’s uncertainty. For applications with a human in the loop, this enables the end user to choose from a small set of plausible candidate orientations, dramatically simplifying the orientation estimation task while preserving user control over the outputs.
Our contributions include the following: (1) we identify fundamental obstacles to orientation estimation and study the conditions under which a naïve regression-based approach to orientation estimation fails; (2) we propose a two-stage orientation estimation pipeline that sidesteps these obstacles; (3) we train and test our model on Shapenet and show that it achieves SOTA performance for orientation estimation; (4) we use conformal prediction to enable end users to resolve ambiguities in a shape’s orientation; (5) we release our code and model weights to share our work with the ML community.
2 Related work
Classical methods.
A simple method for orientation estimation is to compute a rotation that aligns a shape’s principal axes with the coordinate axes; Kaye & Ivrissimtzis (2015) propose a robust variant of this method for mesh alignment. However, Kazhdan et al. (2003) find that PCA-based orientation estimation is not robust to asymmetries. Jin et al. (2012); Wang et al. (2014) propose unsupervised methods that leverage low-rank priors on axis-aligned 2D projections and third-order tensors, respectively, constructed from input shapes. These priors are restrictive, and the resulting orientation pipelines also fail on asymmetric shapes.
Another set of classical methods observe that as many man-made objects are designed to stand on flat surfaces, their up axis is normal to a supporting base. Motivated by this observation, these methods attempt to identify a shape’s supporting base rather than directly infer their up axis. Fu et al. (2008) generate a set of candidate bases, extract geometric features, and combine a random forest and SVM to predict a natural base from the candidates. Lin & Tai (2012) simplify a shape’s convex hull, cluster the resulting facets to obtain a set of candidate bases, and compute a hand-designed score to select the best base. Both of these methods rely heavily on feature engineering and fail on shapes that do not have natural supporting bases.
Deep learning-based methods.
Motivated by the limitations of classical approaches, several works use deep learning for orientation estimation. Liu et al. (2016) train two neural networks on voxel representations of 3D shapes. A first-stage network assigns each shape to one of $C$ classes. Based on this prediction, the shape is routed to one of $C$ second-stage networks that are independently trained to predict the up axis from voxel representations of shapes in their respective classes. This method is unable to handle shapes that lie outside the $C$ classes on which the networks were trained.
Pang et al. (2022) draw inspiration from classical methods and train a segmentation network to predict points that belong to a shape’s supporting base. They fit a plane to the predicted base points and output a normal vector to this plane as the predicted up axis. This method represents the current state of the art for orientation estimation, but struggles to handle shapes without well-defined natural bases and and only predicts a shape’s up axis. In contrast, our method succeeds on general shapes and predicts a full rotation matrix that returns a shape to canonical orientation.
Chen et al. (2021) use reinforcement learning to train a model to gradually rotate a shape into upright orientation. While this algorithm performs well, it is evaluated on few classes and is costly to train. Kim et al. (2020) adopt a similar perspective to Fu et al. (2008), but use ConvNets to extract features for a random forest classifier that predicts a natural base. Poursaeed et al. (2020) use orientation estimation as a pretext task to learn features for shape classification and keypoint prediction.
We also highlight a related literature on canonical alignment. This literature includes works such as Kim et al. (2023); Sajnani et al. (2022); Spezialetti et al. (2020), which seek to map arbitrarily-rotated shapes to a class-consistent pose, as well as Katzir et al. (2022); Sun et al. (2021), which seek to learn pose-invariant representations of 3D shapes. These works only attempt to learn a consistent orientation within each class, but this orientation is not consistent across classes and is not generally aligned with the coordinate axes. In contrast, we tackle the more challenging task of inferring a canonical orientation that is consistent across all objects.
3 Method
In this section, we motivate and describe our orientation pipeline. We first identify fundamental obstacles to orientation estimation and show that learning a shape’s orientation with the $L^{2}$ loss fails when the shape is rotationally symmetric. Motivated by these observations, we introduce our two-stage orientation pipeline consisting of a quotient orienter followed by a flipper. Our quotient orienter model solves a regression problem to recover a shape’s orientation up to octahedral symmetries, which commonly occur in real-world shapes. The flipper then predicts one of 24 octahedral flips that returns the first-stage output to canonical orientation. We finally use conformal prediction to enable our flipper to output prediction sets whose size varies with the model’s uncertainty. This allows end users to resolve ambiguities in a shape’s orientation by choosing from a small set of plausible candidate orientations.
3.1 Orientation estimation under rotational symmetries
Figure 2: A shape’s orientation $\Omega_{S}$ is a rotation matrix whose columns are the shape’s side-, up-, and front-axes (plotted in yellow, magenta, cyan, resp).
In this section, we introduce the orientation estimation problem and motivate our approach. Throughout these preliminaries, we consider 3D shapes $S\in\mathcal{S}$ lying in some space of arbitrary shape representations $\mathcal{S}$. Orientation estimation consists of learning an orienter function $f:\mathcal{S}\rightarrow SO(3)$ that maps a shape $S\in\mathcal{S}$ to a predicted orientation $\hat{\Omega}_{S}\in SO(3)$, where $SO(3)$ denotes the 3D rotation group. An orientation is a rotation matrix $\Omega_{S}$ associated with a shape $S$ that is rotation-equivariant: If one rotates $S$ by $R\in SO(3)$ to obtain $RS$, then $\Omega_{RS}=R\Omega_{S}$. We interpret the columns of $\Omega_{S}=(\omega_{S}^{x},\omega_{S}^{y},\omega_{S}^{z})$ as the side-, up-, and front-axes of $S$, respectively, and say that $S$ is in canonical orientation if $\Omega_{S}=I$. If $S$ is in canonical orientation, then its side-, up-, and front-axes are aligned with the $\{x,y,z\}$ coordinate axes, respectively. We depict a canonically-oriented shape $S$ along with its orientation $\Omega_{S}$ in Figure 2.
Given a training set $\mathcal{D}$ of shapes $S\in\mathcal{S}$ paired with their ground truth orientations $\Omega_{S}$, a natural strategy for orientation estimation is to define a loss function $\ell$ on the space of orientations, parametrize $f$ as a neural network $f_{\theta}$, and solve the following problem:
$\min_{f_{\theta}}\underset{\begin{subarray}{c}R\sim U(SO(3))\\ (S,\Omega_{S})\in\mathcal{D}\end{subarray}}{\mathbb{E}}\left[\ell\left(f_{\theta}(RS),R,\Omega_{S}\right)\right],$ (1)
where $U(SO(3))$ is the uniform distribution over $SO(3)$. In the following section, we will motivate our approach by describing a theoretical obstacle to learning an orienter function $f$ regardless of one’s choice of loss $\ell$, and then describe specific challenges associated with solving Equation 1 using the $L^{2}$ loss.
Theoretical challenges.
An ideal orienter $f$ should satisfy two desiderata: (1) It should accept arbitrarily-oriented shapes $RS$ as input and output their orientation $\Omega_{RS}=R\Omega_{S}$, and (2) it should succeed on most shapes $S$ that occur in the wild. As many real-world shapes possess at least one non-trivial rotational symmetry, an ideal orienter should therefore succeed on rotationally-symmetric shapes. We begin by showing that no function can simultaneously satisfy these desiderata.
Proposition 3.1
Let $S\in\mathcal{S}$ be a fixed shape which is symmetric under a non-trivial group of rotations $\mathcal{R}_{S}\subseteq SO(3)$, and let $\Omega_{S}$ be its orientation. Then there is no function $f$ such that $f(RS)=R\Omega_{S}$ for all $R\in SO(3)$.
We prove this result in Appendix A.1. Intuitively, if a shape $S$ is symmetric under some non-trivial rotation $R$, then $S$ is invariant under $R$, but its orientation is equivariant under $R$, so the map $RS\mapsto R\Omega_{S}$ is one-to-many and cannot be a function. This shows that any solution to the orientation estimation problem 1 necessarily trades off some desirable property; there are no functions that can successfully orient a rotationally-symmetric shape given any input orientation.
Previous works such as Liu et al. (2016); Poursaeed et al. (2020) observe that orientation estimation via $L^{2}$ regression typically yields poor results. Motivated by this observation, we now characterize the solution to orientation estimation via $L^{2}$ regression for a single rotationally-symmetric shape and show that naïve $L^{2}$ regression degenerates in this setting.
Proposition 3.2
Let $S\in\mathcal{S}$ be a fixed shape which is symmetric under a non-trivial group of rotations $\mathcal{R}_{S}\subseteq SO(3)$. Let $\Omega_{S}$ be the shape’s orientation, and suppose $f^{*}:\mathcal{S}\rightarrow SO(3)$ solves the following regression problem:
$\min_{f:\mathcal{S}\rightarrow SO(3)}\underset{R\sim U(SO(3))}{\mathbb{E}}\left[\|f(RS)-R\Omega_{S}\|_{F}^{2}\right],$ (2)
Then $f^{*}(RS)=\textrm{proj}_{SO(3)}\left[\frac{1}{|\mathcal{R}_{S}|}\sum_{Q\in\mathcal{R}_{S}}RQ\Omega_{S}\right]\neq R\Omega_{S}$, where $\textrm{proj}_{SO(3)}$ denotes the orthogonal projection onto $SO(3)$.
We prove this proposition in Appendix A.2. Proposition 3.2 shows that even when seeking to predict the rotated orientations $R\Omega_{S}$ of a single rotationally-symmetric shape $S$, $L^{2}$ regression fails to learn the correct solution, and instead learns the Euclidean mean of the rotated orientations $RQ\Omega_{S}$ across all rotations $Q$ in the symmetry group $\mathcal{R}_{S}$ (Moakher, 2002).
This problem may be highly degenerate, even for shapes with a single non-trivial symmetry. For example, consider the bench shape $S$ depicted in Figures LABEL:fig:teaser, 2, LABEL:fig:prop3.1. As shown in Figure LABEL:fig:prop3.1, this shape has two rotational symmetries: The identity rotation, and a $180^{\circ}$ rotation about the $y$-axis. One may represent these rotations by the matrices $I$ and $Q:=(-e_{x},e_{y},-e_{z})$, resp., where $e_{x},e_{y},e_{z}$ are the standard basis vectors.
Figure 4: The solution $f^{*}(S)$ to Problem 2 evaluated at the bench shape $S$ may be any rotation about the $y$-axis.
Proposition 3.2 states that one solves the $L^{2}$ regression problem 2 for the bench shape by computing the arithmetic mean of $I$ and $Q$ and then orthogonally projecting this matrix onto $SO(3)$. The arithmetic mean of $I,Q$ is the matrix $M:=(0,e_{y},0)$, and one computes its orthogonal projection onto $SO(3)$ by solving a special Procrustes problem (Gower & Dijksterhuis, 2004). However, the solution to this problem is non-unique, and $f^{*}(S)$ may be any rotation about the $y$-axis, which we illustrate in Figure 4. This shows that even a single non-trivial rotational symmetry leads to an entire submanifold of solutions $f^{*}(S)$ to Problem 2.
A partial solution.
The previous section shows that solving Equation 1 with the $L^{2}$ loss fails for rotationally-symmetric shapes, which are common in practice. We now present a partial solution to this problem. Suppose we know a finite group ^$\hat{\mathcal{R}}\supseteq\mathcal{R}_{S}$ that contains a shape $S$’s rotational symmetries. We can then quotient the $L^{2}$ loss by $\hat{\mathcal{R}}$ to obtain the following problem:
$\min_{f:\mathcal{S}\rightarrow SO(3)}\underset{\begin{subarray}{c}R\sim U(SO(3))\\ (S,\Omega_{S})\in\mathcal{D}\end{subarray}}{\mathbb{E}}\left[\min_{Q\in\hat{\mathcal{R}}}\|f(RS)-RQ\Omega_{S}\|_{F}^{2}\right].$ (3)
This loss is small if $f(RS)$ is close to the orientation $RQ\Omega_{S}$ of the rotated shape $RQS$ for any $Q\in\hat{\mathcal{R}}$; Mehr et al. (2018) use similar techniques to learn latent shape representations that are invariant under a group of geometric transformations. Intuitively, whereas Equation 2 attempts to make $f(S)$ close to all $Q\Omega_{S}$, a minimizer of Equation 3 merely needs to make $f(S)$ close to any $Q\Omega_{S}$. Formally:
Proposition 3.3
Let $S\in\mathcal{S}$ be a fixed shape which is symmetric under a group of rotations $\mathcal{R}_{S}\subseteq\hat{\mathcal{R}}\subseteq SO(3)$. Let $\Omega_{S}$ be the shape’s orientation, and suppose $f^{*}:\mathcal{S}\rightarrow SO(3)$ is a solution to the following quotient regression problem:
$\min_{f:\mathcal{S}\rightarrow SO(3)}\underset{R\sim U(SO(3))}{\mathbb{E}}\left[\min_{Q\in\hat{\mathcal{R}}}\|f(RS)-RQ\Omega_{S}\|_{F}^{2}\right],$ (4)
Then for any $R\in SO(3)$, $f^{*}(RS)=RQ^{*}\Omega_{S}$ for some $Q^{*}\in\hat{\mathcal{R}}$.
We prove this proposition in Appendix A.3. In contrast to naïve $L^{2}$ regression, quotient regression learns a function that correctly orients rotationally-symmetric shapes up to a rotation in the group $\hat{\mathcal{R}}$. While this is only a partial solution to the orientation estimation problem, the remainder reduces to a discrete classification problem: Predicting the rotation $Q^{*}\in\hat{\mathcal{R}}$ such that $f^{*}(RS)=RQ^{*}\Omega_{S}$. In the following section, we will show how a solution to this problem allows one to map $RS$ to the canonically-oriented shape $S$.
Recovering an orientation via classification.
By solving the quotient regression problem in Equation 3, one can recover an arbitrarily-rotated shape $RS$’s orientation up to a rotation $Q^{*}\in\hat{\mathcal{R}}$. In this section, we propose training a classifier to predict this rotation $Q^{*}$ given the solution $f^{*}(RS)=RQ^{*}\Omega_{S}$ to the quotient regression problem. We now further assume that the shape $S$’s ground truth orientation $\Omega_{S}$ is the canonical orientation $\Omega_{S}=I$. We show that even if $S$ is symmetric under some group of symmetries $\mathcal{R}_{S}\subseteq\hat{\mathcal{R}}$, the optimal classifier’s predictions enable one to map $RS$ to the canonically-oriented shape $S$.
Predicting a rotation $Q^{*}\in\hat{\mathcal{R}}$ from the output $f^{*}(RS)=RQ^{*}$ of the quotient regression model is an $|\hat{\mathcal{R}}|$-class classification problem. One may train an appropriate classifier by solving the following problem:
$\min_{p_{\phi}:\mathcal{S}\rightarrow\Delta^{|\hat{\mathcal{R}}|-1}}\underset{\begin{subarray}{c}Q\sim U(\hat{\mathcal{R}})\\ S\in\mathcal{D}\end{subarray}}{\mathbb{E}}\left[\textrm{CE}\left(p_{\phi}(QS),\delta_{Q}\right)\right],$ (5)
where $U(\hat{\mathcal{R}})$ denotes the uniform distribution on $\hat{\mathcal{R}}$, $\textrm{CE}(\cdot)$ denotes the cross-entropy loss, and $\delta_{Q}\in\Delta^{|\hat{\mathcal{R}}|-1}$ is a one-hot vector centered at the index of $Q\in\hat{\mathcal{R}}$. While one may hope that composing the quotient regression model and this classifier yields an orienter that outputs correct orientations $\hat{\Omega}_{RS}=R\Omega_{S}$ regardless of its inputs’ symmetries, recall that Proposition 3.1 shows such an orienter does not exist. However, the following result shows that this pipeline recovers the orientation of a shape $RS$ up to one of its rotational symmetries, which is sufficient for mapping $RS$ to $S$.
Proposition 3.4
Let $S\in\mathcal{S}$ be a fixed shape which is symmetric under a group of rotations $\mathcal{R}_{S}\subseteq\hat{\mathcal{R}}\subseteq SO(3)$, and suppose $S$ is canonically-oriented, so $\Omega_{S}=I$. Let $f^{*}:\mathcal{S}\rightarrow SO(3)$ be a solution to Equation 3, so that $f^{*}(RS)=RQ^{*}$ for some $Q^{*}\in\hat{\mathcal{R}}$. Finally, suppose that $p^{*}:\mathcal{S}\rightarrow\Delta^{|\hat{\mathcal{R}}|-1}$ solves the following problem:
$\min_{p:\mathcal{S}\rightarrow\Delta^{|\hat{\mathcal{R}}|-1}}\underset{Q\sim U(\hat{\mathcal{R}})}{\mathbb{E}}\left[\textrm{CE}\left(p(QS),\delta_{Q}\right)\right].$ (6)
Then for any $R\in SO(3)$, $p^{*}(f^{*}(RS)^{\top}RS)$ is the uniform distribution over $\left\{(Q^{*})^{\top}F:F\in\mathcal{R}_{S}\right\}$. For any $(Q^{*})^{\top}F$ in the support of this distribution, $\underbrace{((Q^{*})^{\top}F)^{\top}}_{\textrm{second-stage prediction}}\underbrace{f^{*}(RS)^{\top}}_{\textrm{first-stage prediction}}RS=S$, so using $f^{*}$ and $p^{*}$, one may recover $S$ from the arbitrarily-rotated shape $RS$.
We prove this proposition in Appendix A.4. How does one reconcile this result with Proposition 3.1? The orientation of $(Q^{*})^{\top}F)^{\top}f^{*}(RS)^{\top}RS$ is $F^{\top}\neq I=\Omega_{S}$, so this result does not contradict Proposition 3.1. Rather, it shows that when a shape $S$ is rotationally symmetric, one need only predict the orientation of $RS$ up to one of its symmetries to recover the canonically-oriented $S$. We combine these results in the following section to implement a state-of-the-art method for orientation estimation.
3.2 Implementation
Informed by our insights from Section 3.1, we now present our state-of-the-art method for orientation estimation. Our pipeline consists of two components. Our first component, which we call the quotient orienter, is a neural network trained to solve Problem 3. We quotient the $L^{2}$ objective by ^$\hat{\mathcal{R}}:=\mathcal{O}\subseteq SO(3)$, the octahedral group containing the 24 rotational symmetries of a cube. This is among the largest finite subgroups of $SO(3)$ (only the cyclic group $C_{n}$ for $n\geq 48$ and dihedral group $D_{n}$ for $n\geq 4$ can contain more subgroups), and it includes many rotational symmetries that commonly occur in real-world shapes. Our second component, which we call the flipper, is a neural network trained to solve the classification problem 5. We illustrate the output of each stage of this pipeline in Figure LABEL:fig:our-pipeline-bench. As many shapes possess multiple plausible orientations, we use conformal prediction to enable our flipper to output adaptive prediction sets whose size varies with the flipper model’s uncertainty. We provide further implementation details below.
Quotient orienter.
We parametrize our quotient orienter by a DGCNN (Wang et al., 2019) operating on point clouds. To ensure that our predicted orientations lie in $SO(3)$, we follow Brégier (2021) and map model outputs from $\mathbb{R}^{3\times 3}$ to $SO(3)$ by solving the special orthogonal Procrustes problem. We train the quotient orienter on point clouds sampled from the surfaces of meshes in Shapenet (Chang et al., 2015). As these meshes are pre-aligned to lie in canonical orientation, we fix $\Omega_{S}=I$ for all training shapes $S$. We provide full architecture and training details in Appendix B.
In our experiments, we observe that our quotient orienter yields accurate predictions for most input rotations $R$ but fails for a small subset of rotations. To handle this, we follow Liu et al. (2016) and employ test-time augmentation to improve our model’s predictions. This consists of (1) randomly rotating the inputs $RS$ by $K$ random rotations $R_{k}\sim U(SO(3))$, $k=1,...,K$, (2) obtaining the quotient orienter’s predictions $f_{\theta}(R_{k}RS)$ for each shape, (3) returning these predictions to the original input’s orientation by computing $R_{k}^{\top}f_{\theta}(R_{k}RS)$, and (4) outputting the prediction $R_{k^{*}}^{\top}f_{\theta}(R_{k^{*}}RS)$ with the smallest average quotient distance to the remaining predictions.
Flipper.
We also parametrize our flipper by a DGCNN operating on point clouds. We train the flipper on point clouds sampled from the surface of Shapenet meshes by optimizing Equation 5. We draw rotations $Q\sim U(\mathcal{O})$ during training, and simulate inaccuracies in our quotient orienter’s predictions by further rotating the training shapes about a randomly drawn axis by an angle uniformly drawn from $[0,10]$ degrees. We provide full architecture and training details in Appendix B.
We also employ test-time augmentation to improve our flipper model’s predictions. Similarly to the case with the quotient orienter, we (1) randomly flip the inputs by $K$ random rotations $R_{k}\sim\hat{\mathcal{R}}=\mathcal{O}$, (2) obtain the flipper’s predictions for each shape, (3) return these predictions to the original input’s orientation, and (4) output the plurality prediction.
Adaptive prediction sets.
Many real-world shapes have several plausible canonical orientations, even when they lack rotational symmetries. Furthermore, the flipper model may map nearly-symmetric shapes with unique canonical orientations to a uniform distribution over their near-symmetries due to factors such as insufficiently dense point clouds or the smoothness of the flipper function.
To mitigate this issue in pipelines with a human in the loop, we enable our flipper model to output adaptive prediction sets whose size varies with the flipper’s uncertainty (Romano et al., 2020). This method uses a small conformal calibration set drawn from the validation data to learn a tuning parameter $\tau>0$ that controls the size of the prediction sets. Given the flipper model’s output probabilities $p_{\phi}(S)\in\Delta^{|\hat{\mathcal{R}}|-1}$ for some shape $S$, one sorts $p_{\phi}(S)$ in descending order and adds elements of $\hat{\mathcal{R}}$ to the prediction set until their total mass in $p_{\phi}(S)$ reaches $\tau$. Intuitively, these sets will be small when the flipper is confident in its prediction and assigns large mass to the highest-probability classes. Conversely, the sets will be large when the flipper is uncertain and assigns similar mass to most classes.
4 Experiments
We now evaluate our method’s performance on orientation estimation. We first follow the evaluation procedure in Pang et al. (2022) and benchmark against their “Upright-Net,” which represents the current state of the art for orientation estimation. Upright-Net can only map shapes into upright orientation, where a shape’s up axis is aligned with the $y$-axis; in contrast, our method recovers a full orientation $\Omega_{S}$ for each shape. We therefore follow this benchmark with an evaluation of our method on the more challenging task of full-orientation estimation. We incorporate adaptive prediction sets at this stage and demonstrate that our method reliably provides a plausible set of candidate orientations for diverse shapes unseen during training.
4.1 Upright orientation estimation
We construct a random 90-10 train-test split of Shapenet, draw 10k point samples from the surface of each mesh, and train our quotient orienter and flipper on all classes in the training split. We train our quotient orienter for 1919 epochs and our flipper for 3719 epochs, sampling 2k points per point cloud at each iteration and fixing a learning rate of $10^{-4}$. We also train Upright-Net with 2048 points per cloud on the same data for 969 epochs at the same learning rate, at which point the validation accuracy has plateaued. We follow the annotation procedure in Pang et al. (2022) to obtain ground truth segmentations of each point cloud into supporting base points and non-base points.
We then follow the evaluation procedure in Pang et al. (2022) to benchmark our method against their SOTA method for upright orientation estimation. We randomly rotate shapes $S$ in the validation set, use our two-stage pipeline and Upright-Net to estimate the up-axis $\omega_{RS}^{y}$ of each randomly rotated shape $RS$, and then measure the angular error $\textrm{arccos}(\langle\hat{\omega}_{RS}^{y},\omega_{RS}^{y}\rangle)$ between the estimated and ground truth up-axis. Our method’s estimated up-axis is the second column of our estimated orientation matrix $\hat{\Omega}_{RS}$. In contrast, Upright-Net predicts a set of base points for $RS$, fits a plane to these points, and returns this plane’s normal vector pointing towards the shape’s center of mass. We follow Pang et al. (2022) and define our methods’ respective accuracies to be the proportion of validation meshes whose angular error is less than $10^{\circ}$.
Table 1: Up-axis estimation accuracy for our pipeline trained on Shapenet
Method Accuracy ($\uparrow$)
Shapenet ModelNet40
Ours 89.2 % 77.7 %
Upright-Net 69.5 % 62.3 %
We depict the results of this benchmark in Table 1. Our method improves on Upright-Net’s up-axis estimation accuracy by nearly 20 percentage points, corresponding to a 64.6% reduction in the error rate relative to the previous state of the art. To provide a more comprehensive picture of our respective models’ performance, we also report angular loss histograms for our model and Upright-Net in the left panel of Figure LABEL:fig:up-axis-loss-histograms. Our model primarily fails by outputting orientations that are $90^{\circ}$ or $180^{\circ}$ away from the correct orientation, which correspond to failures of the flipper. In contrast, Upright-Net’s failures are more evenly distributed across angular errors. Finally, we depict a grid of non-cherry-picked outputs of our model and Upright-Net in Figure LABEL:fig:sec-4.1-grid and highlight each model’s failure cases in red.
We quantitatively evaluate our model’s generalization by performing the same experiment on ModelNet40 (Wu et al., 2015). Both models’ performances deteriorate in this setting, but our algorithm continues to substantially outperform Upright-Net. Furthermore, the right panel of Figure LABEL:fig:up-axis-loss-histograms shows that our model’s failures on ModelNet40 are more heavily weighted towards flipper failures (where the angular error is close to $90^{\circ}$ and $180^{\circ}$). In the following section, we will show how a human in the loop can resolve these failures by choosing from a small set of candidate flips, which substantially improves our pipeline’s quantitative performance.
These results demonstrate that our method significantly improves over the state of the art in upright orientation estimation. In the following section, we show that our method also successfully recovers the full orientation $\Omega_{RS}$ of a rotated shape, a more challenging task than the well-studied task of upright orientation estimation. Using our estimated orientations, we return a wide variety of shapes into canonical orientation.
4.2 Full-orientation estimation
We now evaluate our method’s performance on full-orientation estimation, in which we use our model’s full orientation matrix $\hat{\Omega}_{RS}$ to transform an arbitrarily-rotated shape $RS$ to the canonically-oriented shape $S$. To our knowledge, our algorithm is the first to solve this task for generic shapes without requiring class information at training time or at inference time. We now record the angular distance between our estimated orientations $\hat{\Omega}_{RS}$ and the ground truth orientations $\Omega_{RS}=R\Omega_{S}=R$. This angular distance is defined as $d(\hat{\Omega}_{RS},R):=\textrm{arccos}(\frac{\textrm{tr}R_{\textrm{diff}}-1}{2})$, where $R_{\textrm{diff}}:=\hat{\Omega}_{RS}R^{\top}$.
As noted in Section 3.2, many real-world shapes have several plausible canonical orientations, and our flipper may also map nearly-symmetric shapes to a uniform distribution over their near-symmetries. In particular, while most real-world shapes have a well-defined upright orientation, their front-facing orientation is often ambiguous. To account for this, we incorporate adaptive prediction sets at this stage of our evaluation. We measure the angular distance between the estimated orientations $\hat{\Omega}^{k}_{RS}$ corresponding to the top $K=4$ flips in our flipper’s model distribution and the ground truth orientation $\Omega_{RS}$, and take their maximum to obtain our reported angular errors. This resolves ambiguities in a shape’s front-facing direction (there are 4 possible front-facing directions for each upright orientation) and also simulates the ability of a human in the loop to choose between a small set of candidate orientations.
We describe our method’s performance on full-orientation estimation in Table 2, where we adopt a $10^{\circ}$ accuracy threshold for full rotation estimation. Our method achieves high top-4 accuracy on full-orientation estimation, but its accuracy deteriorates in the $K=1$ case, where one can only consider the flipper’s highest-probability class. This is partially attributable to ambiguities in the front-facing orientation of many shapes. To demonstrate this, the left panel of Figure LABEL:fig:sec-4.2-grid depicts shapes where our method’s top-1 angular error is $>10^{\circ}$ in red. These failures primarily correspond to shapes that are symmetric with respect to rotations about the $y$-axis; these shapes are correctly oriented even though our method has recovered incorrect orientations.
Table 2: Full-orientation estimation accuracy
Method Accuracy ($\uparrow$)
Top-1 Top-4
Ours 68.3 % 85.8 %
Figure LABEL:fig:sec-4.2-grid compares the shapes obtained using our model’s outputs when $K=1$ (where we only consider the flipper’s highest-probability predicted flip) and when $K=4$ (where we depict the best flip among the top 4 classes in the flipper’s model distribution) to the canonically-oriented shapes. We highlight the model’s failure cases in red. Even in the $K=1$ case, many of the model’s failures correspond to orientations that are plausible or correct up to a symmetry of the shape. We bolster this claim with additional non-cherry-picked examples in Appendix C; see Figures 10, 11.
Finally, in Figure 12 in Appendix C, we depict transformed shapes obtained by applying our method to randomly-rotated shapes from the Objaverse dataset (Deitke et al., 2023). This dataset contains highly diverse meshes of varying quality and therefore serves as a useful test case for our method’s performance on out-of-distribution shapes. (As these meshes are not consistently oriented in the dataset, we cannot train on them or report meaningful error metrics.) Using our orientation pipeline, one reliably recovers shapes that are canonically-oriented up to an octahedral flip. Our flipper has greater difficulty handling out-of-distribution meshes, but predicts an acceptable flip in many cases. We expect that training our flipper on a larger dataset of oriented shapes will further improve its generalization performance.
5 Conclusion
This work introduces a state-of-the-art method for 3D orientation estimation. Whereas previous approaches can only infer upright orientations for limited classes of shapes, our method successfully recovers entire orientations for general shapes. We show that naïve regression-based approaches for orientation estimation degenerate on rotationally-symmetric shapes, which are common in practice, and develop a two-stage orientation pipeline that avoids these obstacles. Our pipeline first orients an arbitrarily rotated input shape up to an octahedral symmetry, and then predicts the octahedral symmetry that maps the first-stage output to the canonically-oriented shape.
In the language of group theory, our two-stage approach first learns to map an input shape $RS$ onto the orbit of the desired output $S$ under some symmetry group $\hat{\mathcal{R}}$, and then learns a distribution over elements of this orbit. We anticipate that this factorization of geometric learning problems will be broadly applicable throughout 3D deep learning for tackling problems that are ill-posed due to the presence of symmetries. We also believe that our results can be further improved by training our quotient orienter and flipper models on larger datasets of consistently-oriented shapes as they become available.
Acknowledgments
We thank the authors of Pang et al. (2022) for their gracious assistance with their source code. CS thanks Backflip AI for hosting him and generously supporting him with mentorship and compute throughout his internship.
Appendix A Proofs
A.1 Proposition 3.1
Suppose $f(RS)=R\Omega_{S}$ for all $R\in SO(3)$, and fix some non-identity rotation $R\in\mathcal{R}_{S}$ under which $S$ is symmetric. Then $RS=S$, but $f(RS)=R\Omega_{S}\neq\Omega_{S}=f(S)$, so $f(RS)\neq f(S)$ even though $RS=S$. Hence $f$ must be a one-to-many map and is therefore not a function. $\blacksquare$
A.2 Proposition 3.2
The key insight is that if $f$ is a function, then $f(RS)=f(R^{\prime}S)$ for all $R^{\prime}\in SO(3)$ such that $RS=R^{\prime}S$. Equation 2 will then drive the optimal $f^{*}(RS)$ to the Euclidean mean (Moakher, 2002) of the rotation matrices $R^{\prime}$ such that $RS=R^{\prime}S$. We begin by showing that these are precisely the matrices $RQ$ for $Q\in\mathcal{R}_{S}$.
As $\mathcal{R}_{S}$ is the group of symmetries of $S$, $QS=S$ for all $Q\in\mathcal{R}_{S}$. Given some rotation $R\in SO(3)$, left-multiplying by $R$ then yields $RQS=RS$ for all $Q\in\mathcal{R}_{S}$. This relationship also holds in reverse: If $RS=R^{\prime}S$ for $R,R^{\prime}\in SO(3)$, then $R^{\prime}=RQ$ for some $Q\in\mathcal{R}_{S}$. To see this, note that if $RS=R^{\prime}S$, then $S=R^{\top}R^{\prime}S$ and hence $R^{\top}R^{\prime}\in\mathcal{R}_{S}$. Consequently, $R^{\prime}=R(R^{\top}R^{\prime})=RQ$ for $Q:=R^{\top}R^{\prime}\in\mathcal{R}_{S}$. It follows that:
$\left\{R^{\prime}\in SO(3):RS=R^{\prime}S\right\}=\left\{RQ:Q\in\mathcal{R}_{S}\right\}.$
We can therefore write a solution to Equation 2 evaluated at $RS$ as follows:
$f^{*}(RS)$ $=\underset{R^{*}\in SO(3)}{\textrm{argmin}}\underset{R^{\prime}\in SO(3):RS=R^{\prime}S}{\mathbb{E}}\left[\|R^{*}-R^{\prime}\Omega_{S}\|_{F}^{2}\right]$
$=\underset{R^{*}\in SO(3)}{\textrm{argmin}}\underset{RQ:Q\in\mathcal{R}_{S}}{\mathbb{E}}\left[\|R^{*}-RQ\Omega_{S}\|_{F}^{2}\right]$
$=\underset{R^{*}\in SO(3)}{\textrm{argmin}}\underset{Q\in U(\mathcal{R}_{S})}{\mathbb{E}}\left[\|R^{*}-RQ\Omega_{S}\|_{F}^{2}\right]$
$=\underset{R^{*}\in SO(3)}{\textrm{argmin}}\frac{1}{|\mathcal{R}_{S}|}\sum_{Q\in\mathcal{R}_{S}}\|R^{*}-RQ\Omega_{S}\|_{F}^{2}.$
This is the Euclidean mean of the matrices $RQ\Omega_{S}$ as defined in Moakher (2002). Proposition 3.3 in the same reference states that the solution to this problem is found by computing the arithmetic mean $\frac{1}{|\mathcal{R}_{S}|}\sum_{Q\in\mathcal{R}_{S}}RQ\Omega_{S}$ and then orthogonally projecting this onto $SO(3)$. In particular,
$f^{*}(RS)=\textrm{proj}_{SO(3)}\left[\frac{1}{|\mathcal{R}_{S}|}\sum_{Q\in\mathcal{R}_{S}}RQ\Omega_{S}\right]\neq R\Omega_{S}.$
Hence $L^{2}$ regression fails to learn the orientation $\Omega_{S}$ of a shape $S\in\mathcal{S}$ that possesses a non-trivial set of rotational symmetries $\mathcal{R}_{S}$. $\blacksquare$
A.3 Proposition 3.3
We begin by defining an equivalence relation over $SO(3)$. Given two rotations $R_{1},R_{2}\in SO(3)$, we call $R_{1},R_{2}$ equivalent and write $R_{1}\sim R_{2}$ if there exists some $Q\in\hat{\mathcal{R}}$ such that $R_{2}=R_{1}Q$. We verify that this is an equivalence relation:
Reflexivity:
$I\in\hat{\mathcal{R}}$ since $\hat{\mathcal{R}}$ is a group and $R_{1}=R_{1}I$, so $R_{1}\sim R_{1}$.
Symmetry:
Suppose $R_{1}\sim R_{2}$. Then $R_{2}=R_{1}Q$ for some $Q\in\hat{\mathcal{R}}$. As $\hat{\mathcal{R}}$ is a group, $R^{\top}=R^{-1}\in\hat{\mathcal{R}}$ as well, and $R_{2}Q^{\top}=R_{1}$, so $R_{2}\sim R_{1}$.
Transitivity:
Suppose $R_{1}\sim R_{2}$ and $R_{2}\sim R_{3}$. Then there are $Q,Q^{\prime}\in\hat{\mathcal{R}}$ such that $R_{2}=R_{1}Q$ and $R_{3}=R_{2}Q^{\prime}$. Hence $R_{3}=R_{2}Q^{\prime}=R_{1}QQ^{\prime}$, and as $\hat{\mathcal{R}}$ is a group, $QQ^{\prime}\in\hat{\mathcal{R}}$. We conclude that $R_{1}\sim R_{3}$.
This confirms that $\sim$ is a valid equivalence relation. Using this equivalence relation, we partition $SO(3)$ into equivalence classes, choose a unique representative for each class, and use $[R]\in SO(3)/\sim$ to denote the unique representative for the equivalence class containing $R\in SO(3)$. We then use this map to define a candidate solution to Equation 4 over the space of rotated shapes $\{RS:R\in SO(3)\}$ as $f^{*}(RS):=[R]\Omega_{S}$. We will first verify that this defines a valid function (i.e. that $f^{*}$ is not one-to-many), and then show that it attains a loss value of 0 in Equation 4.
We first show that $f^{*}$ defines a valid function. To do so, we must show that if $R_{1}S=R_{2}S$, then $f^{*}(R_{1}S)=f^{*}(R_{2}S)$. To this end, suppose that $R_{1}S=R_{2}S$. Then $S=R_{1}^{\top}R_{2}S$, so $Q:=R_{1}^{\top}R_{2}\in\mathcal{R}_{S}\subseteq\hat{\mathcal{R}}$. It follows that $R_{2}=R_{1}R_{1}^{\top}R_{2}=R_{1}Q$ for some $Q\in\hat{\mathcal{R}}$, so $R_{1}\sim R_{2}$. Since $R_{1}\sim R_{2}$, $[R_{1}]=[R_{2}]$ and so $f^{*}(R_{1}S)=[R_{1}]\Omega_{S}=[R_{2}]\Omega_{S}=f^{*}(R_{2}S)$. This shows that $f^{*}$ defines a valid function.
We now show that $f^{*}$ attains a loss value of 0 in Equation 4. For any $R\in SO(3)$, we have:
$\min_{Q\in\hat{\mathcal{R}}}\|f(RS)-RQ\Omega_{S}\|_{F}^{2}=\min_{Q\in\hat{\mathcal{R}}}\|[R]\Omega_{S}-RQ\Omega_{S}\|_{F}^{2}.$
But clearly $R\sim[R]$, so there exists some $Q^{*}\in\hat{\mathcal{R}}$ such that $[R]=RQ^{*}$. Hence
$\min_{Q\in\hat{\mathcal{R}}}\|[R]\Omega_{S}-RQ\Omega_{S}\|_{F}^{2}=0,$
and as this reasoning holds for any $R\in SO(3)$, it follows that
$\underset{R\sim U(SO(3))}{\mathbb{E}}\left[\min_{Q\in\hat{\mathcal{R}}}\|f(RS)-RQ\Omega_{s}\|_{F}^{2}\right]=0.$
We conclude that $f^{*}$ is a minimizer of Equation 4. Furthermore, $f^{*}(S)=[I]\Omega_{S}=Q^{*}\Omega_{S}$ for some $Q^{*}\in\hat{\mathcal{R}}$, which completes the proof of the proposition. $\blacksquare$
A.4 Proposition 3.4
If $F\in\mathcal{R}_{S}$, then $FS=S$, so $QFS=QS$ for any other rotation $Q\in SO(3)$ and $\{QFS:F\in\mathcal{R}_{S}\}$ contains the symmetries of the rotated shape $QS$. The optimal solution $p^{*}$ to Equation 6 maps a rotated shape $QS$ (where $Q\in\hat{\mathcal{R}}$) to the empirical distribution of the targets $Q\in\hat{\mathcal{R}}$ conditional on a shape $QS$. But if $QFS=QS$ for all $F\in\mathcal{R}_{S}$, then this is the uniform distribution over the set $\left\{QF:F\in\mathcal{R}_{S}\right\}$.
Since $f^{*}(RS)=RQ^{*}$ for some $Q^{*}\in\hat{\mathcal{R}}$, $f^{*}(RS)^{\top}RS=(Q^{*})^{\top}S$, and applying the general result from above, we conclude that $p^{*}(f^{*}(RS)^{\top}RS)=p^{*}((Q^{*})^{\top}S)$ is the uniform distribution over the set $\{(Q^{*})^{\top}F:F\in\mathcal{R}_{S}\}$.
For any $(Q^{*})^{\top}F$, one then computes $((Q^{*})^{\top}F)^{\top}f^{*}(RS)^{\top}RS=F^{\top}S$. But as $\mathcal{R}_{S}$ is a group, $F^{\top}\in\mathcal{R}_{S}$ whenever $F$ is, so $F^{\top}S=S$ and we conclude that $((Q^{*})^{\top}F)^{\top}f^{*}(RS)^{\top}RS=S$. $\blacksquare$
Appendix B Implementation details
B.1 Quotient orienter
We parametrize our quotient orienter by a DGCNN and use the author’s Pytorch implementation (Wang et al., 2019) with 1024-dimensional embeddings, $k=20$ neighbors for the EdgeConv layers, and a dropout probability of 0.5. Our DGCNN outputs unstructured $3\times 3$ matrices, which we then project onto $SO(3)$ by solving a special orthogonal Procrustes problem; we use the roma package (Brégier, 2021) to efficiently compute this projection.
We train our quotient orienter on point clouds consisting of 10k surface samples from Shapenet meshes. We subsample 2k points per training iteration and pass batches of 48 point clouds per iteration. We train the quotient orienter for 1919 epochs at a learning rate of $10^{-4}$.
For test-time augmentation, we (1) randomly rotate the inputs $RS$ by $K$ random rotations $R_{k}\sim U(SO(3))$, $k=1,...,K$, (2) obtain the quotient orienter’s predictions $f_{\theta}(R_{k}RS)$ for each shape, (3) return these predictions to the original input’s orientation by computing $R_{k}^{\top}f_{\theta}(R_{k}RS)$, and (4) output the prediction $R_{k^{*}}^{\top}f_{\theta}(R_{k^{*}}RS)$ with the smallest average quotient distance to the remaining predictions.
B.2 Flipper
We parametrize our flipper by a DGCNN and use the author’s Pytorch implementation (Wang et al., 2019) with 1024-dimensional embeddings, $k=20$ neighbors for the EdgeConv layers, and a dropout probability of 0.5. Our flipper outputs $24$-dimensional logits, as we quotient our first-stage regression problem by the octahedral group, which contains the 24 rotational symmetries of a cube.
We train our flipper on point clouds consisting of 10k surface samples from Shapenet meshes. We subsample 2k points per training iteration and pass batches of 48 point clouds per iteration. We train the quotient orienter for 3719 epochs at a learning rate of $10^{-4}$. We draw rotations $Q\in U(\mathcal{O})$ during training, and simulate inaccuracies in our quotient orienter’s predictions by further rotating the training shapes about a randomly drawn axis by an angle uniformly drawn from $[0,10]$ degrees.
We also employ test-time augmentation to improve our flipper model’s predictions. Similarly to the case with the quotient orienter, we (1) randomly flip the inputs by $K$ random rotations $R_{k}\sim\hat{\mathcal{R}}=\mathcal{O}$, (2) obtain the flipper’s predictions for each shape, (3) return these predictions to the original input’s orientation, and (4) output the plurality prediction.
B.3 Adaptive prediction sets
We implement adaptive prediction sets following the method in Angelopoulos et al. (2020) with their regularization parameter $\lambda$ set to 0. To calibrate our conformal flipper, we first draw a subset of the validation set (the calibration set), apply a random octahedral flip $Q\sim U(\mathcal{O})$ to each calibration shape, and then pass each flipped shape $QS$ through the trained flipper to obtain class probabilities $p_{\phi}(QS)\in\Delta^{2}3$. The calibration score for a shape $S$ is the sum of the model’s class probabilities $p(QS)_{i}$ ranked in descending order, up to and including the true class $i^{*}$ corresponding to the ground truth flip $Q$. We fix a confidence level $1-\alpha$ and return the $(1-\alpha)$-th quantile $\tau$ of the calibration scores for each shape in the calibration set. In general, smaller values of $1-\alpha$ lead to smaller values of $\tau$, which ultimately results in smaller prediction sets at inference time, whereas large values of $1-\alpha$ lead to larger prediction sets at inference time but with stronger guarantees that these sets include the true flip.
At inference time, we first obtain the flipper model’s output probabilities $p_{\phi}(S)\in\Delta^{|\hat{\mathcal{R}}|-1}$ for some shape $S$, then sort $p_{\phi}(S)$ in descending order and add elements of $\hat{\mathcal{R}}$ to the prediction set until their total mass in $p_{\phi}(S)$ reaches $\tau$. Intuitively, these sets will be small when the flipper is confident in its prediction and assigns large mass to the highest-probability classes. Conversely, the sets will be large when the flipper is uncertain and assigns similar mass to most classes.
Appendix C Additional results
Figure 10: A grid of our model’s top-1 outputs given randomly rotated, non-cherrypicked meshes from Shapenet. We depict a front view of our model’s recovered shapes along with an inset depicting an isometric view and highlight failure cases in red. Even when our model fails to recover the correct orientation, the recovered shape is often acceptable in practice.
Figure 11: A grid of our model’s top-4 outputs given randomly rotated, non-cherrypicked meshes from Shapenet. We depict a front view of our model’s recovered shapes along with an inset depicting an isometric view and highlight failure cases in red.
Figure 12: A grid of our model’s top-1 outputs given randomly rotated, non-cherrypicked meshes from Objaverse. We depict a front view of our model’s recovered shapes along with an inset depicting an isometric view. Our quotient regressor consistently succeeds on out of distribution meshes, as most of our pipeline’s outputs are correctly oriented up to a cube flip. Our flipper has greater difficulty generalizing, but predicts an acceptable flip in many cases. We expect that training on a larger and more diverse dataset of oriented shapes will improve our flipper’s generalization performance.

### v2

@@
 We motivate our engineering contributions by theory describing fundamental obstacles to orientation estimation for rotationally-symmetric shapes, and show how our method avoids these obstacles.
+3d orientation, shape analysis, 3d deep learning, geometric deep learning
 1 Introduction
@@
 Using this data, one can rotate a shape into canonical orientation, in which the shape’s orientation axes are aligned with the coordinate axes.
-This task is especially important as a pre-processing step in 3D deep learning, where deep networks are typically trained on datasets of canonically-oriented shapes but applied to arbitrarily-oriented shapes at inference time.
+This task is especially important as a pre-processing step in 3D deep learning, where deep networks are typically trained on datasets of canonically oriented shapes but applied to arbitrarily-oriented shapes at inference time.
 While data augmentation or equivariant and invariant architectures may improve a model’s robustness to input rotations, these techniques come at the cost of data efficiency and model expressivity (Kuchnik & Smith, 2019; Kim et al., 2023).
 In contrast, orientation estimation allows one to pre-process shapes at inference time so that their orientation matches a model’s training data.
-Orientation estimation is a challenging task, and developing an orientation pipeline that reliably estimates complete orientations of general shapes remains an open problem.
-The naïve deep learning approach is to train a model with an $L^{2}$ loss to directly predict a shape’s orientation from a point cloud of surface samples.
-However, this strategy fails for shapes with rotational symmetries, where the optimal solution to the $L^{2}$ regression problem is the Euclidean mean (Moakher, 2002) of a shape’s orientations over all of its symmetries.
+A shape’s orientation axes are not intrinsic geometric quantities: They are determined by humans based on physical and functional considerations.
+It is therefore difficult to construct simple geometric algorithms for orientation estimation.
+However, given a large dataset of canonically oriented shapes, one may pose orientation estimation as a supervised learning problem.
+This task is challenging, and developing an orientation pipeline that reliably estimates complete orientations of general shapes remains an open problem.
+The naïve deep learning approach is to train a model with an $L_{2}$ loss to directly predict a shape’s orientation from a point cloud of surface samples.
+However, this strategy fails for shapes with rotational symmetries, where the optimal solution to the $L_{2}$ regression problem is the Euclidean mean (Moakher, 2002) of a shape’s orientations over all of its symmetries.
 In contrast, works such as Poursaeed et al.
@@
 Consequently, quotienting our first-stage regression problem by octahedral symmetries prevents its predictions from collapsing to averages, while also keeping the subsequent classification problem tractable.
-Using this strategy, our method achieves state-of-the-art performance on the well-studied problem of up-axis prediction, and additionally performs well on full-orientation prediction, which few prior works have tackled.
+Using this strategy, our method accurately estimates a shape’s orientation up to one of its symmetries, which suffices for returning the shape to an upright and front-facing pose.
+As we are unaware of prior work on this problem, we also benchmark our pipeline’s performance on the well-studied task of up-axis prediction, in which one seeks to return a shape to an upright (but not necessarily front-facing) pose, and find that our work achieves state-of-the-art performance on this task.
 Unlike previous work, we train and evaluate our model on the entire Shapenet dataset rather than a subset of classes.
@@
 For applications with a human in the loop, this enables the end user to choose from a small set of plausible candidate orientations, dramatically simplifying the orientation estimation task while preserving user control over the outputs.
-Our contributions include the following: (1) we identify fundamental obstacles to orientation estimation and study the conditions under which a naïve regression-based approach to orientation estimation fails; (2) we propose a two-stage orientation estimation pipeline that sidesteps these obstacles; (3) we train and test our model on Shapenet and show that it achieves SOTA performance for orientation estimation; (4) we use conformal prediction to enable end users to resolve ambiguities in a shape’s orientation; (5) we release our code and model weights to share our work with the ML community.
+Our contributions include the following: (1) we identify fundamental obstacles to orientation estimation and study the conditions under which a naïve regression-based approach to orientation estimation fails; (2) we propose a two-stage orientation estimation pipeline that sidesteps these obstacles; (3) we train and test our model on Shapenet and show that it achieves SOTA performance for up-axis prediction and orientation estimation; (4) we use conformal prediction to enable end users to resolve ambiguities in a shape’s orientation; (5) we release our code and model weights to share our work with the ML community.
 2 Related work
@@
 (2020) use orientation estimation as a pretext task to learn features for shape classification and keypoint prediction.
+They also investigate a pure classification-based approach to orientation estimation that discretizes the 3D rotation group into $K$ rotations and predicts a distribution over these rotations for an arbitrarily-rotated input shape.
+They find that its performance decays rapidly as $K$ increases, reaching an accuracy as low as 1.6% for $K=100$ rotations.
 We also highlight a related literature on canonical alignment.
@@
 (2022); Spezialetti et al.
+(2020); Zhou et al.
 (2022), which seek to map arbitrarily-rotated shapes to a class-consistent pose, as well as Katzir et al.
@@
 In contrast, we tackle the more challenging task of inferring a canonical orientation that is consistent across all objects.
+More broadly, geometric deep learning studies general methods for exploiting symmetries in data by designing machine learning models that are invariant or equivariant to certain transformations.
+For instance, Kaba et al.
+(2023) proposes using learned canonicalization functions to obtain equivariant machine learning methods, and Puny et al.
+(2022) proposes frame averaging to induce invariance or equivariance in deep learning architectures.
+Our work may be viewed as an efficient canonicalization method for the specific case of 3D shapes with rotational symmetries.
 3 Method
 In this section, we motivate and describe our orientation pipeline.
-We first identify fundamental obstacles to orientation estimation and show that learning a shape’s orientation with the $L^{2}$ loss fails when the shape is rotationally symmetric.
+We first identify fundamental obstacles to orientation estimation and show that learning a shape’s orientation with the $L_{2}$ loss fails when the shape is rotationally symmetric.
 Motivated by these observations, we introduce our two-stage orientation pipeline consisting of a quotient orienter followed by a flipper.
@@
 3.1 Orientation estimation under rotational symmetries
-Figure 2: A shape’s orientation $\Omega_{S}$ is a rotation matrix whose columns are the shape’s side-, up-, and front-axes (plotted in yellow, magenta, cyan, resp).
 In this section, we introduce the orientation estimation problem and motivate our approach.
@@
 We interpret the columns of $\Omega_{S}=(\omega_{S}^{x},\omega_{S}^{y},\omega_{S}^{z})$ as the side-, up-, and front-axes of $S$, respectively, and say that $S$ is in canonical orientation if $\Omega_{S}=I$.
-If $S$ is in canonical orientation, then its side-, up-, and front-axes are aligned with the $\{x,y,z\}$ coordinate axes, respectively.
-We depict a canonically-oriented shape $S$ along with its orientation $\Omega_{S}$ in Figure 2.
-Given a training set $\mathcal{D}$ of shapes $S\in\mathcal{S}$ paired with their ground truth orientations $\Omega_{S}$, a natural strategy for orientation estimation is to define a loss function $\ell$ on the space of orientations, parametrize $f$ as a neural network $f_{\theta}$, and solve the following problem:
-$\min_{f_{\theta}}\underset{\begin{subarray}{c}R\sim U(SO(3))\\ (S,\Omega_{S})\in\mathcal{D}\end{subarray}}{\mathbb{E}}\left[\ell\left(f_{\theta}(RS),R,\Omega_{S}\right)\right],$ (1)
+If $S$ is in canonical orientation, then its side-, up-, and front-axes (which we will jointly refer to as its orientation axes) are aligned with the $\{x,y,z\}$ coordinate axes, respectively.
+We depict a canonically oriented shape $S$ along with its orientation $\Omega_{S}$ in Figure 2.
+Figure 2: A shape’s orientation $\Omega_{S}$ is a rotation matrix whose columns are the shape’s side-, up-, and front-axes (plotted in yellow, magenta, cyan, resp).
+A shape’s orientation axes are not intrinsic geometric quantities.
+Instead, they are determined by humans based on physical and functional considerations.
+For instance, a shape’s up-axis may be determined by how a shape naturally lies under gravity, whereas its front-axis may be determined by the direction from which a human typically interacts with it.
+It is therefore challenging to devise simple geometric algorithms for orientation estimation.
+However, there exist large datasets of shapes such as Shapenet (Chang et al., 2015) that have been assigned a canonical orientation by human annotators.
+Given a training set $\mathcal{D}$ of shapes $S\in\mathcal{S}$ paired with their ground truth orientations $\Omega_{S}$ (where $\Omega_{S}\equiv I$ for datasets of canonically oriented shapes), a natural strategy for orientation estimation is to parametrize the orienter as a neural network $f_{\theta}$ and solve the following supervised learning problem:
+$\min_{f_{\theta}}\underset{\begin{subarray}{c}R\sim U(SO(3))\\ (S,\Omega_{S})\in\mathcal{D}\end{subarray}}{\mathbb{E}}\left[\|f_{\theta}(RS)-R\Omega_{S}\|_{F}^{2}\right],$ (1)
 where $U(SO(3))$ is the uniform distribution over $SO(3)$.
-In the following section, we will motivate our approach by describing a theoretical obstacle to learning an orienter function $f$ regardless of one’s choice of loss $\ell$, and then describe specific challenges associated with solving Equation 1 using the $L^{2}$ loss.
-Theoretical challenges.
-An ideal orienter $f$ should satisfy two desiderata: (1) It should accept arbitrarily-oriented shapes $RS$ as input and output their orientation $\Omega_{RS}=R\Omega_{S}$, and (2) it should succeed on most shapes $S$ that occur in the wild.
-As many real-world shapes possess at least one non-trivial rotational symmetry, an ideal orienter should therefore succeed on rotationally-symmetric shapes.
-We begin by showing that no function can simultaneously satisfy these desiderata.
-Proposition 3.1
-Let $S\in\mathcal{S}$ be a fixed shape which is symmetric under a non-trivial group of rotations $\mathcal{R}_{S}\subseteq SO(3)$, and let $\Omega_{S}$ be its orientation.
-Then there is no function $f$ such that $f(RS)=R\Omega_{S}$ for all $R\in SO(3)$.
-We prove this result in Appendix A.1.
-Intuitively, if a shape $S$ is symmetric under some non-trivial rotation $R$, then $S$ is invariant under $R$, but its orientation is equivariant under $R$, so the map $RS\mapsto R\Omega_{S}$ is one-to-many and cannot be a function.
-This shows that any solution to the orientation estimation problem 1 necessarily trades off some desirable property; there are no functions that can successfully orient a rotationally-symmetric shape given any input orientation.
-Previous works such as Liu et al.
-(2016); Poursaeed et al.
-(2020) observe that orientation estimation via $L^{2}$ regression typically yields poor results.
-Motivated by this observation, we now characterize the solution to orientation estimation via $L^{2}$ regression for a single rotationally-symmetric shape and show that naïve $L^{2}$ regression degenerates in this setting.
-Proposition 3.2
+Many real-world shapes $S$ possess at least one non-trivial rotational symmetry $R$.
+This rotation leaves the shape unchanged, so $RS=S$.
+However, since a shape’s orientation is rotation-equivariant, $\Omega_{RS}=R\Omega_{S}\neq\Omega_{S}$, and the same shape $S$ will necessarily be associated with two distinct orientations $\Omega_{S}$ and $R\Omega_{S}$.
+It follows that the ground truth orienter map $S\mapsto\Omega_{S}$ of a rotationally-symmetric shape is one-to-many and therefore not a function.
+We formally state and prove this result in Proposition A.1, and depict an instance of this phenomenon in Figure LABEL:fig:prop3.1.
+On the other hand, $f_{\theta}$ is a function, which implies that the solution to Equation 1 cannot be the true orienter map.
+The following proposition shows that for single shapes with rotational symmetries, the solution to Equation 1 is instead Euclidean mean of the rotated orientations $RQ\Omega_{S}$ across all rotations $Q$ in the symmetry group $\mathcal{R}_{S}$ (Moakher, 2002).
+Proposition 3.1.
 Let $S\in\mathcal{S}$ be a fixed shape which is symmetric under a non-trivial group of rotations $\mathcal{R}_{S}\subseteq SO(3)$.
@@
 We prove this proposition in Appendix A.2.
-Proposition 3.2 shows that even when seeking to predict the rotated orientations $R\Omega_{S}$ of a single rotationally-symmetric shape $S$, $L^{2}$ regression fails to learn the correct solution, and instead learns the Euclidean mean of the rotated orientations $RQ\Omega_{S}$ across all rotations $Q$ in the symmetry group $\mathcal{R}_{S}$ (Moakher, 2002).
 This problem may be highly degenerate, even for shapes with a single non-trivial symmetry.
@@
 As shown in Figure LABEL:fig:prop3.1, this shape has two rotational symmetries: The identity rotation, and a $180^{\circ}$ rotation about the $y$-axis.
-One may represent these rotations by the matrices $I$ and $Q:=(-e_{x},e_{y},-e_{z})$, resp., where $e_{x},e_{y},e_{z}$ are the standard basis vectors.
+It is straightforward to show that the solution $f^{*}$ to Equation 2 is non-unique and may be any rotation about the $y$-axis, which we illustrate in Figure 4.
+(See Appendix A.3 for further details.) This demonstrates that even a single non-trivial rotational symmetry leads to an entire submanifold of solutions $f^{*}(S)$ to Problem 2, which substantiates the observations of Liu et al.
+(2016); Poursaeed et al.
+(2020) that orientation estimation via $L_{2}$ regression typically performs poorly in practice.
 Figure 4: The solution $f^{*}(S)$ to Problem 2 evaluated at the bench shape $S$ may be any rotation about the $y$-axis.
-Proposition 3.2 states that one solves the $L^{2}$ regression problem 2 for the bench shape by computing the arithmetic mean of $I$ and $Q$ and then orthogonally projecting this matrix onto $SO(3)$.
-The arithmetic mean of $I,Q$ is the matrix $M:=(0,e_{y},0)$, and one computes its orthogonal projection onto $SO(3)$ by solving a special Procrustes problem (Gower & Dijksterhuis, 2004).
-However, the solution to this problem is non-unique, and $f^{*}(S)$ may be any rotation about the $y$-axis, which we illustrate in Figure 4.
-This shows that even a single non-trivial rotational symmetry leads to an entire submanifold of solutions $f^{*}(S)$ to Problem 2.
-A partial solution.
-The previous section shows that solving Equation 1 with the $L^{2}$ loss fails for rotationally-symmetric shapes, which are common in practice.
+3.2 A partial solution.
+The previous section shows that orientation estimation via $L_{2}$ regression fails for rotationally-symmetric shapes, which are common in practice.
 We now present a partial solution to this problem.
 Suppose we know a finite group ^$\hat{\mathcal{R}}\supseteq\mathcal{R}_{S}$ that contains a shape $S$’s rotational symmetries.
-We can then quotient the $L^{2}$ loss by $\hat{\mathcal{R}}$ to obtain the following problem:
+We can then quotient the $L_{2}$ loss by $\hat{\mathcal{R}}$ to obtain the following problem:
 $\min_{f:\mathcal{S}\rightarrow SO(3)}\underset{\begin{subarray}{c}R\sim U(SO(3))\\ (S,\Omega_{S})\in\mathcal{D}\end{subarray}}{\mathbb{E}}\left[\min_{Q\in\hat{\mathcal{R}}}\|f(RS)-RQ\Omega_{S}\|_{F}^{2}\right].$ (3)
@@
 Formally:
-Proposition 3.3
+Proposition 3.2.
 Let $S\in\mathcal{S}$ be a fixed shape which is symmetric under a group of rotations $\mathcal{R}_{S}\subseteq\hat{\mathcal{R}}\subseteq SO(3)$.
@@
 We prove this proposition in Appendix A.4.
-In contrast to naïve $L^{2}$ regression, quotient regression learns a function that correctly orients rotationally-symmetric shapes up to a rotation in the group $\hat{\mathcal{R}}$.
+In contrast to naïve $L_{2}$ regression, quotient regression learns a function that correctly orients rotationally-symmetric shapes up to a rotation in the group $\hat{\mathcal{R}}$.
 While this is only a partial solution to the orientation estimation problem, the remainder reduces to a discrete classification problem: Predicting the rotation $Q^{*}\in\hat{\mathcal{R}}$ such that $f^{*}(RS)=RQ^{*}\Omega_{S}$.
-In the following section, we will show how a solution to this problem allows one to map $RS$ to the canonically-oriented shape $S$.
-Recovering an orientation via classification.
+In the following section, we will show how a solution to this problem allows one to map $RS$ to the canonically oriented shape $S$.
+3.3 Recovering an orientation via classification.
 By solving the quotient regression problem in Equation 3, one can recover an arbitrarily-rotated shape $RS$’s orientation up to a rotation $Q^{*}\in\hat{\mathcal{R}}$.
@@
 We now further assume that the shape $S$’s ground truth orientation $\Omega_{S}$ is the canonical orientation $\Omega_{S}=I$.
-We show that even if $S$ is symmetric under some group of symmetries $\mathcal{R}_{S}\subseteq\hat{\mathcal{R}}$, the optimal classifier’s predictions enable one to map $RS$ to the canonically-oriented shape $S$.
+We show that even if $S$ is symmetric under some group of symmetries $\mathcal{R}_{S}\subseteq\hat{\mathcal{R}}$, the optimal classifier’s predictions enable one to map $RS$ to the canonically oriented shape $S$.
 Predicting a rotation $Q^{*}\in\hat{\mathcal{R}}$ from the output $f^{*}(RS)=RQ^{*}$ of the quotient regression model is an $|\hat{\mathcal{R}}|$-class classification problem.
-One may train an appropriate classifier by solving the following problem:
-$\min_{p_{\phi}:\mathcal{S}\rightarrow\Delta^{|\hat{\mathcal{R}}|-1}}\underset{\begin{subarray}{c}Q\sim U(\hat{\mathcal{R}})\\ S\in\mathcal{D}\end{subarray}}{\mathbb{E}}\left[\textrm{CE}\left(p_{\phi}(QS),\delta_{Q}\right)\right],$ (5)
-where $U(\hat{\mathcal{R}})$ denotes the uniform distribution on $\hat{\mathcal{R}}$, $\textrm{CE}(\cdot)$ denotes the cross-entropy loss, and $\delta_{Q}\in\Delta^{|\hat{\mathcal{R}}|-1}$ is a one-hot vector centered at the index of $Q\in\hat{\mathcal{R}}$.
-While one may hope that composing the quotient regression model and this classifier yields an orienter that outputs correct orientations $\hat{\Omega}_{RS}=R\Omega_{S}$ regardless of its inputs’ symmetries, recall that Proposition 3.1 shows such an orienter does not exist.
-However, the following result shows that this pipeline recovers the orientation of a shape $RS$ up to one of its rotational symmetries, which is sufficient for mapping $RS$ to $S$.
-Proposition 3.4
-Let $S\in\mathcal{S}$ be a fixed shape which is symmetric under a group of rotations $\mathcal{R}_{S}\subseteq\hat{\mathcal{R}}\subseteq SO(3)$, and suppose $S$ is canonically-oriented, so $\Omega_{S}=I$.
+While one may hope that composing the quotient regression model and the optimal classifier yields a function that outputs correct orientations $\hat{\Omega}_{RS}=R\Omega_{S}$ regardless of its inputs’ symmetries, such a function cannot exist for symmetric shapes.
+However, the following result shows that this pipeline recovers the orientation of a shape $RS$ up to one of its rotational symmetries, which suffices for mapping $RS$ to $S$.
+Proposition 3.3.
+Let $S\in\mathcal{S}$ be a fixed shape which is symmetric under a group of rotations $\mathcal{R}_{S}\subseteq\hat{\mathcal{R}}\subseteq SO(3)$, and suppose $S$ is canonically oriented, so $\Omega_{S}=I$.
 Let $f^{*}:\mathcal{S}\rightarrow SO(3)$ be a solution to Equation 3, so that $f^{*}(RS)=RQ^{*}$ for some $Q^{*}\in\hat{\mathcal{R}}$.
 Finally, suppose that $p^{*}:\mathcal{S}\rightarrow\Delta^{|\hat{\mathcal{R}}|-1}$ solves the following problem:
-$\min_{p:\mathcal{S}\rightarrow\Delta^{|\hat{\mathcal{R}}|-1}}\underset{Q\sim U(\hat{\mathcal{R}})}{\mathbb{E}}\left[\textrm{CE}\left(p(QS),\delta_{Q}\right)\right].$ (6)
+$\min_{p:\mathcal{S}\rightarrow\Delta^{|\hat{\mathcal{R}}|-1}}\underset{Q\sim U(\hat{\mathcal{R}})}{\mathbb{E}}\left[\textrm{CE}\left(p(QS),\delta_{Q}\right)\right],$ (5)
+where $U(\hat{\mathcal{R}})$ denotes the uniform distribution on $\hat{\mathcal{R}}$, $\textrm{CE}(\cdot)$ denotes the cross-entropy loss, and $\delta_{Q}\in\Delta^{|\hat{\mathcal{R}}|-1}$ is a one-hot vector centered at the index of $Q\in\hat{\mathcal{R}}$.
 Then for any $R\in SO(3)$, $p^{*}(f^{*}(RS)^{\top}RS)$ is the uniform distribution over $\left\{(Q^{*})^{\top}F:F\in\mathcal{R}_{S}\right\}$.
@@
 We prove this proposition in Appendix A.5.
-How does one reconcile this result with Proposition 3.1?
-The orientation of $(Q^{*})^{\top}F)^{\top}f^{*}(RS)^{\top}RS$ is $F^{\top}\neq I=\Omega_{S}$, so this result does not contradict Proposition 3.1.
-Rather, it shows that when a shape $S$ is rotationally symmetric, one need only predict the orientation of $RS$ up to one of its symmetries to recover the canonically-oriented $S$.
+The results in this section show that unlike naïve $L_{2}$ regression, a two-stage pipeline consisting of quotient regression followed by discrete classification can successfully recover a shape’s orientation up to its symmetries, provided one quotients the $L_{2}$ objective by a sufficiently large subgroup of $SO(3)$.
 We combine these results in the following section to implement a state-of-the-art method for orientation estimation.
 3.4 Implementation
-Informed by our insights from Section 3.1, we now present our state-of-the-art method for orientation estimation.
+Informed by our insights from Sections 3.2 and 3.3, we now present our state-of-the-art method for orientation estimation.
 Our pipeline consists of two components.
 Our first component, which we call the quotient orienter, is a neural network trained to solve Problem 3.
-We quotient the $L^{2}$ objective by ^$\hat{\mathcal{R}}:=\mathcal{O}\subseteq SO(3)$, the octahedral group containing the 24 rotational symmetries of a cube.
+We quotient the $L_{2}$ objective by ^$\hat{\mathcal{R}}:=\mathcal{O}\subseteq SO(3)$, the octahedral group containing the 24 rotational symmetries of a cube.
 This is among the largest finite subgroups of $SO(3)$ (only the cyclic group $C_{n}$ for $n\geq 48$ and dihedral group $D_{n}$ for $n\geq 4$ can contain more subgroups), and it includes many rotational symmetries that commonly occur in real-world shapes.
-Our second component, which we call the flipper, is a neural network trained to solve the classification problem 5.
+Our second component, which we call the flipper, is a classifier trained to predict the rotation $Q^{*}\in\hat{\mathcal{R}}$ from the output $f^{*}(RS)=RQ^{*}$ of the quotient regression model.
 We illustrate the output of each stage of this pipeline in Figure LABEL:fig:our-pipeline-bench.
@@
 We also parametrize our flipper by a DGCNN operating on point clouds.
-We train the flipper on point clouds sampled from the surface of Shapenet meshes by optimizing Equation 5.
+We train the flipper on point clouds sampled from the surface of Shapenet meshes by solving the $|\hat{\mathcal{R}}|$-class classification problem described in Section 3.3.
 We draw rotations $Q\sim U(\mathcal{O})$ during training, and simulate inaccuracies in our quotient orienter’s predictions by further rotating the training shapes about a randomly drawn axis by an angle uniformly drawn from $[0,10]$ degrees.
@@
 To mitigate this issue in pipelines with a human in the loop, we enable our flipper model to output adaptive prediction sets whose size varies with the flipper’s uncertainty (Romano et al., 2020).
-This method uses a small conformal calibration set drawn from the validation data to learn a tuning parameter $\tau>0$ that controls the size of the prediction sets.
+This method uses a small conformal calibration set drawn from the validation data to learn a threshold parameter $\tau>0$ associated with a coverage probability $\alpha\in(0,1)$ that controls the size of the prediction sets.
 Given the flipper model’s output probabilities $p_{\phi}(S)\in\Delta^{|\hat{\mathcal{R}}|-1}$ for some shape $S$, one sorts $p_{\phi}(S)$ in descending order and adds elements of $\hat{\mathcal{R}}$ to the prediction set until their total mass in $p_{\phi}(S)$ reaches $\tau$.
@@
 (2022) and benchmark against their “Upright-Net,” which represents the current state of the art for orientation estimation.
-Upright-Net can only map shapes into upright orientation, where a shape’s up axis is aligned with the $y$-axis; in contrast, our method recovers a full orientation $\Omega_{S}$ for each shape.
+Upright-Net can only map shapes into upright orientation, where a shape’s up-axis is aligned with the $y$-axis; in contrast, our method recovers a full orientation $\Omega_{S}$ for each shape.
 We therefore follow this benchmark with an evaluation of our method on the more challenging task of full-orientation estimation.
 We incorporate adaptive prediction sets at this stage and demonstrate that our method reliably provides a plausible set of candidate orientations for diverse shapes unseen during training.
+We train and evaluate all models on Shapenet (Chang et al., 2015), as this is the largest and most diverse dataset we are aware of containing canonically oriented shapes.
+However, we report qualitative results for our method’s out-of-distribution performance on Objaverse in Appendix C.
 4.1 Upright orientation estimation
@@
 (2022) to obtain ground truth segmentations of each point cloud into supporting base points and non-base points.
+Table 1: Up-axis estimation accuracy for our pipeline trained on Shapenet
+Method Accuracy ($\uparrow$)
+Shapenet ModelNet40
+Ours 89.2 % 77.7 %
+Upright-Net (Pang et al., 2022) 69.5 % 62.3 %
 We then follow the evaluation procedure in Pang et al.
@@
 In contrast, Upright-Net predicts a set of base points for $RS$, fits a plane to these points, and returns this plane’s normal vector pointing towards the shape’s center of mass.
+This method relies on a restrictive prior on the geometry of the input shapes that does not hold for shapes which do not naturally lie on a supporting base.
 We follow Pang et al.
 (2022) and define our methods’ respective accuracies to be the proportion of validation meshes whose angular error is less than $10^{\circ}$.
-Table 1: Up-axis estimation accuracy for our pipeline trained on Shapenet
-Method Accuracy ($\uparrow$)
-Shapenet ModelNet40
-Ours 89.2 % 77.7 %
-Upright-Net 69.5 % 62.3 %
 We depict the results of this benchmark in Table 1.
 Our method improves on Upright-Net’s up-axis estimation accuracy by nearly 20 percentage points, corresponding to a 64.6% reduction in the error rate relative to the previous state of the art.
-To provide a more comprehensive picture of our respective models’ performance, we also report angular loss histograms for our model and Upright-Net in the left panel of Figure LABEL:fig:up-axis-loss-histograms.
+To provide a more comprehensive picture of our respective models’ performance, we also report the empirical CDF of angular errors for our model and Upright-Net in the left panel of Figure LABEL:fig:up-axis-loss-histograms.
 Our model primarily fails by outputting orientations that are $90^{\circ}$ or $180^{\circ}$ away from the correct orientation, which correspond to failures of the flipper.
 In contrast, Upright-Net’s failures are more evenly distributed across angular errors.
-Finally, we depict a grid of non-cherry-picked outputs of our model and Upright-Net in Figure LABEL:fig:sec-4.1-grid and highlight each model’s failure cases in red.
+Finally, we depict a grid of non-cherry-picked outputs of our model and Upright-Net in Figure LABEL:fig:sec-4.1-grid and highlight failure cases in red.
 We quantitatively evaluate our model’s generalization by performing the same experiment on ModelNet40 (Wu et al., 2015).
@@
 In the following section, we will show how a human in the loop can resolve these failures by choosing from a small set of candidate flips, which substantially improves our pipeline’s quantitative performance.
-These results demonstrate that our method significantly improves over the state of the art in upright orientation estimation.
-In the following section, we show that our method also successfully recovers the full orientation $\Omega_{RS}$ of a rotated shape, a more challenging task than the well-studied task of upright orientation estimation.
+These results demonstrate that our method significantly improves over the state of the art in up-axis estimation.
+In the following section, we show that our method also successfully recovers the full orientation $\Omega_{RS}$ of a rotated shape up to its symmetries, a more challenging task than upright orientation estimation.
 Using our estimated orientations, we return a wide variety of shapes into canonical orientation.
 4.2 Full-orientation estimation
-We now evaluate our method’s performance on full-orientation estimation, in which we use our model’s full orientation matrix $\hat{\Omega}_{RS}$ to transform an arbitrarily-rotated shape $RS$ to the canonically-oriented shape $S$.
+We now evaluate our method’s performance on full-orientation estimation, in which we use our model’s full predicted orientation matrix $\hat{\Omega}_{RS}$ to transform an arbitrarily-rotated shape $RS$ to the canonically oriented shape $S$.
+We draw shapes $S$ from the Shapenet validation set and randomly rotate them to obtain inputs $RS$ to our pipeline.
+Because several orientation matrices may be associated with a single canonically oriented shape when this shape has rotational symmetries, we report the symmetric chamfer distance between our pipeline’s output shape and the ground truth shape.
+Because this is a shape-to-shape metric, it is insensitive to rotational symmetries: If $RS=S$, then $d(S,T)=d(RS,T)$, where $d$ is the chamfer distance and $T$ is a reference shape.
 To our knowledge, our algorithm is the first to solve this task for generic shapes without requiring class information at training time or at inference time.
-We now record the angular distance between our estimated orientations $\hat{\Omega}_{RS}$ and the ground truth orientations $\Omega_{RS}=R\Omega_{S}=R$.
-This angular distance is defined as $d(\hat{\Omega}_{RS},R):=\textrm{arccos}(\frac{\textrm{tr}R_{\textrm{diff}}-1}{2})$, where $R_{\textrm{diff}}:=\hat{\Omega}_{RS}R^{\top}$.
-As noted in Section 3.2, many real-world shapes have several plausible canonical orientations, and our flipper may also map nearly-symmetric shapes to a uniform distribution over their near-symmetries.
-In particular, while most real-world shapes have a well-defined upright orientation, their front-facing orientation is often ambiguous.
-To account for this, we incorporate adaptive prediction sets at this stage of our evaluation.
-We measure the angular distance between the estimated orientations $\hat{\Omega}^{k}_{RS}$ corresponding to the top $K=4$ flips in our flipper’s model distribution and the ground truth orientation $\Omega_{RS}$, and take their maximum to obtain our reported angular errors.
-This resolves ambiguities in a shape’s front-facing direction (there are 4 possible front-facing directions for each upright orientation) and also simulates the ability of a human in the loop to choose between a small set of candidate orientations.
-We describe our method’s performance on full-orientation estimation in Table 2, where we adopt a $10^{\circ}$ accuracy threshold for full rotation estimation.
-Our method achieves high top-4 accuracy on full-orientation estimation, but its accuracy deteriorates in the $K=1$ case, where one can only consider the flipper’s highest-probability class.
-This is partially attributable to ambiguities in the front-facing orientation of many shapes.
-To demonstrate this, the left panel of Figure LABEL:fig:sec-4.2-grid depicts shapes where our method’s top-1 angular error is $>10^{\circ}$ in red.
-These failures primarily correspond to shapes that are symmetric with respect to rotations about the $y$-axis; these shapes are correctly oriented even though our method has recovered incorrect orientations.
-Table 2: Full-orientation estimation accuracy
-Method Accuracy ($\uparrow$)
-Top-1 Top-4
-Ours 68.3 % 85.8 %
-Figure LABEL:fig:sec-4.2-grid compares the shapes obtained using our model’s outputs when $K=1$ (where we only consider the flipper’s highest-probability predicted flip) and when $K=4$ (where we depict the best flip among the top 4 classes in the flipper’s model distribution) to the canonically-oriented shapes.
-We highlight the model’s failure cases in red.
-Even in the $K=1$ case, many of the model’s failures correspond to orientations that are plausible or correct up to a symmetry of the shape.
-We bolster this claim with additional non-cherry-picked examples in Appendix C; see Figures 10, 11.
+We therefore construct two baselines using Upright-Net.
+As this algorithm only predicts a shape’s up-axis (the second column of its orientation matrix), we must augment Upright-Net’s predictions with a front-axis and a side-axis.
+We generate a baseline called “Upright-Net-Oracle” by augmenting Upright-Net’s predicted up-axis with the shape’s ground truth front-axis, and a second baseline called “Upright-Net-Random” by augmenting the predicted up-axis with a random vector that we orthogonalize with respect to the up-axis.
+In both cases, we obtain the side-axis by taking the cross product of the respective up- and front-axes.
+“Upright-Net-Oracle” upper-bounds the performance of a hypothetical orienter built on Upright-Net, whereas “Upright-Net-Random” provides a lower bound on such an orienter’s performance.
+We benchmark our pipeline’s performance (“Orient Anything”) against each baseline in Table 2, where we report the mean chamfer distance between each model’s predicted shapes and the corresponding ground truth shapes, which are in canonical orientation.
+Orient Anything achieves an 84% reduction in mean chamfer distance between predicted and ground truth shapes relative to the oracle baseline, which augments Upright-Net’s predicted up-axis with the shape’s ground truth front-axis.
+In practice, one does not have access to this front-axis at inference time, so this represents a loose upper bound on the potential performance of a shape orientation algorithm built upon Upright-Net’s predicted up-axes.
+This empirically validates our Proposition 3.2 and demonstrates that our pipeline successfully recovers canonically oriented shapes, even when symmetries preclude the orienter map from being represented by a function.
+We further illustrate our pipeline’s performance in Figure 9 in Appendix C, where we depict a large grid of shapes that have been oriented by our pipeline.
+Even when our model’s outputs disagree with the ground truth pose in Shapenet (we render the corresponding meshes in red), the recovered shape is often acceptable in practice.
 Finally, in Figure 11 in Appendix C, we depict transformed shapes obtained by applying our method to randomly-rotated shapes from the Objaverse dataset (Deitke et al., 2023).
 This dataset contains highly diverse meshes of varying quality and therefore serves as a useful test case for our method’s performance on out-of-distribution shapes.
-(As these meshes are not consistently oriented in the dataset, we cannot train on them or report meaningful error metrics.) Using our orientation pipeline, one reliably recovers shapes that are canonically-oriented up to an octahedral flip.
+(As these shapes are not canonically oriented, we cannot train on them or report meaningful error metrics.) Using our orientation pipeline, one reliably recovers shapes that are canonically oriented up to an octahedral flip.
 Our flipper has greater difficulty handling out-of-distribution meshes, but predicts an acceptable flip in many cases.
 We expect that training our flipper on a larger dataset of oriented shapes will further improve its generalization performance.
+Table 2: Full-orientation estimation performance on the Shapenet validation set.
+Method Mean chamfer distance ($\downarrow$)
+Upright-Net-Random 0.10801 $\pm$ 0.13824
+Upright-Net-Oracle 0.05481 $\pm$ 0.13016
+Orient Anything 0.00856 $\pm$ 0.0396
+Orient Anything-APS 0.00208 $\pm$ 0.01407
+Figure 8: Histogram of APS sizes with coverage probability $\alpha=0.3$. The median APS contains 2 shapes, and 90% of APSs contain at most 8 shapes.
+Adaptive prediction sets.
+Many real-world shapes have several plausible canonical orientations, even if they lack rotational symmetries.
+In particular, while most real-world shapes have a well-defined upright orientation, their front-facing orientation is often ambiguous.
+In Section 3.4, we proposed enabling a human in the loop to resolve these ambiguities by having our orientation pipeline output adaptive prediction sets (APSs), whose size varies with the flipper’s uncertainty.
+We now incorporate this technique into our evaluation.
+We learn a threshold parameter $\tau$ corresponding to a coverage probability of $\alpha=0.3$, and at inference time output a set of flips in descending order of probability under the flipper’s output distribution until their total probability mass reaches $\tau$.
+Intuitively, this set will be small when the flipper is confident in its prediction and large otherwise.
+By applying each flip to the first-stage orienter’s output, we obtain a shape orientation pipeline that outputs sets of candidate shapes $\hat{S}_{c}$ given a single arbitrarily-rotated input $RS$.
+We call this method “Orient Anything-APS.” To evaluate its performance, we compute the chamfer distance between each candidate shape and the ground truth shape, and take the minimum over these chamfer distances.
+As one may trivially reduce the minimum chamfer distance over an APS by outputting arbitrarily large sets, we report a histogram of APS sizes in Figure 8 to demonstrate that these sets are typically small.
+We report the performance of “Orient Anything-APS” in the final row of Table 2.
+Allowing our pipeline to output an APS reduces the mean chamfer distance between the closest shape in this set to the ground truth by a factor of 4 relative to Orient Anything without this feature.
+Figure 8 shows that with our chosen coverage probability, most APSs are small; the median APS size is 2, and 90% of APSs have at most 8 shapes.
+This demonstrates that our flipper typically assigns most of its probability mass to a few octahedral symmetries, which often contain a rotation that returns the shape to its canonical pose in Shapenet.
+In Figure 10 in Appendix C, we also include a grid of meshes in the prediction sets associated with 20 randomly-selected validation shapes.
+These prediction sets typically consist of a small collection of plausible shape orientations; by manually inspecting them, a human in the loop may select the orientation that is most suitable for their target application.
 5 Conclusion
@@
 We show that naïve regression-based approaches for orientation estimation degenerate on rotationally-symmetric shapes, which are common in practice, and develop a two-stage orientation pipeline that avoids these obstacles.
-Our pipeline first orients an arbitrarily rotated input shape up to an octahedral symmetry, and then predicts the octahedral symmetry that maps the first-stage output to the canonically-oriented shape.
-In the language of group theory, our two-stage approach first learns to map an input shape $RS$ onto the orbit of the desired output $S$ under some symmetry group $\hat{\mathcal{R}}$, and then learns a distribution over elements of this orbit.
+Our pipeline first orients an arbitrarily rotated input shape up to an octahedral symmetry, and then predicts the octahedral symmetry that maps the first-stage output to the canonically oriented shape.
 We anticipate that this factorization of geometric learning problems will be broadly applicable throughout 3D deep learning for tackling problems that are ill-posed due to the presence of symmetries.
-We also believe that our results can be further improved by training our quotient orienter and flipper models on larger datasets of consistently-oriented shapes as they become available.
-Acknowledgments
+Impact statement.
+This paper presents a novel algorithm for 3D shape orientation. We do not believe that there are notable ethical considerations or societal impacts of our work meriting specific consideration.
+Acknowledgments.
 We thank the authors of Pang et al.
-(2022) for their gracious assistance with their source code.
+(2022) for their gracious assistance with their source code and Justin Solomon for valuable discussions.
 CS thanks Backflip AI for hosting him and generously supporting him with mentorship and compute throughout his internship.
-Appendix A Proofs
-A.1 Proposition 3.1
-Suppose $f(RS)=R\Omega_{S}$ for all $R\in SO(3)$, and fix some non-identity rotation $R\in\mathcal{R}_{S}$ under which $S$ is symmetric.
+Appendix A Additional results and proofs
+A.1 The orienter map for symmetric shapes is not a function
+Proposition A.1.
+Let $S\in\mathcal{S}$ be a fixed shape which is symmetric under a non-trivial group of rotations $\mathcal{R}_{S}\subseteq SO(3)$, and let $\Omega_{S}$ be its orientation.
+Then there is no function $f$ such that $f(RS)=R\Omega_{S}$ for all $R\in SO(3)$.
+Proof: Suppose $f(RS)=R\Omega_{S}$ for all $R\in SO(3)$, and fix some non-identity rotation $R\in\mathcal{R}_{S}$ under which $S$ is symmetric.
 Then $RS=S$, but $f(RS)=R\Omega_{S}\neq\Omega_{S}=f(S)$, so $f(RS)\neq f(S)$ even though $RS=S$.
@@
 $f^{*}(RS)=\textrm{proj}_{SO(3)}\left[\frac{1}{|\mathcal{R}_{S}|}\sum_{Q\in\mathcal{R}_{S}}RQ\Omega_{S}\right]\neq R\Omega_{S}.$
-Hence $L^{2}$ regression fails to learn the orientation $\Omega_{S}$ of a shape $S\in\mathcal{S}$ that possesses a non-trivial set of rotational symmetries $\mathcal{R}_{S}$. $\blacksquare$
-A.3 Proposition 3.3
+Hence $L_{2}$ regression fails to learn the orientation $\Omega_{S}$ of a shape $S\in\mathcal{S}$ that possesses a non-trivial set of rotational symmetries $\mathcal{R}_{S}$. $\blacksquare$
+A.3 Non-uniqueness of solution to Problem 2
+Consider the bench shape $S$ depicted in Figures LABEL:fig:teaser, 2, LABEL:fig:prop3.1.
+As shown in Figure LABEL:fig:prop3.1, this shape has two rotational symmetries: The identity rotation, and a $180^{\circ}$ rotation about the $y$-axis.
+One may represent these rotations by the matrices $I$ and $Q:=(-e_{x},e_{y},-e_{z})$, respectively, where $e_{x},e_{y},e_{z}$ are the standard basis vectors.
+Proposition 3.1 states that one solves the $L^{2}$ regression problem 2 for the bench shape by computing the arithmetic mean of $I$ and $Q$ and then orthogonally projecting this matrix onto $SO(3)$.
+The arithmetic mean of $I,Q$ is the matrix $M:=(0,e_{y},0)$, and one computes its orthogonal projection onto $SO(3)$ by solving a special Procrustes problem (Gower & Dijksterhuis, 2004):
+$\min_{R\in SO(3)}\|R-M\|_{F}^{2}$ (6)
+The solution to this problem is non-unique for $M:=(0,e_{y},0)$, and the minimum is attained by any rotation about the $y$-axis (i.e.
+any rotation matrix whose second solumn is $e_{y}$).
+This shows that even a single non-trivial rotational symmetry leads to an entire submanifold of solutions $f^{*}(S)$ to Problem 2.
+A.4 Proposition 3.2
 We begin by defining an equivalence relation over $SO(3)$.
@@
 $\blacksquare$
-A.4 Proposition 3.4
+A.5 Proposition 3.3
 If $F\in\mathcal{R}_{S}$, then $FS=S$, so $QFS=QS$ for any other rotation $Q\in SO(3)$ and $\{QFS:F\in\mathcal{R}_{S}\}$ contains the symmetries of the rotated shape $QS$.
@@
 Conversely, the sets will be large when the flipper is uncertain and assigns similar mass to most classes.
-Appendix C Additional results
-Figure 10: A grid of our model’s top-1 outputs given randomly rotated, non-cherrypicked meshes from Shapenet.
-We depict a front view of our model’s recovered shapes along with an inset depicting an isometric view and highlight failure cases in red.
-Even when our model fails to recover the correct orientation, the recovered shape is often acceptable in practice.
-Figure 11: A grid of our model’s top-4 outputs given randomly rotated, non-cherrypicked meshes from Shapenet.
-We depict a front view of our model’s recovered shapes along with an inset depicting an isometric view and highlight failure cases in red.
-Figure 12: A grid of our model’s top-1 outputs given randomly rotated, non-cherrypicked meshes from Objaverse.
-We depict a front view of our model’s recovered shapes along with an inset depicting an isometric view.
+Appendix C Additional figures
+Figure 9: A grid of our model’s outputs given randomly rotated, non-cherrypicked meshes from Shapenet.
+We render each mesh whose chamfer distance with respect to the ground truth is below a predetermined success threshold in blue, and we render the remainder in red.
+Our success threshold is the average chamfer distance between a shape randomly rotated by $10^{\circ}$ and the original shape; we compute this average across 100 Shapenet validation meshes.
+Even when predicted shapes do not agree with the ground truth orientation determined by the creators of Shapenet, their orientation is often acceptable in practice.
+Figure 10: A grid of 20 adaptive prediction sets generated using our pipeline given randomly-rotated, non-cherrypicked meshes from Shapenet.
+We render each mesh whose chamfer distance with respect to the ground truth is below a predetermined success threshold in blue, and we render the remainder in red.
+Our success threshold is the average chamfer distance between a shape randomly rotated by $10^{\circ}$ and the original shape; we compute this average across 100 Shapenet validation meshes.
+Our adaptive prediction sets are typically small and include reasonable orientations, which an end user may choose between as needed for their target application.
+Figure 11: A grid of our model’s outputs given randomly rotated, non-cherrypicked meshes from Objaverse.
 Our quotient regressor consistently succeeds on out of distribution meshes, as most of our pipeline’s outputs are correctly oriented up to a cube flip.

### v3

@@
-Orient Anything
+Symmetry-Robust 3D Orientation Estimation
 Abstract
@@
 Let $S\in\mathcal{S}$ be a fixed shape which is symmetric under a group of rotations $\mathcal{R}_{S}\subseteq\hat{\mathcal{R}}\subseteq SO(3)$.
-Let $\Omega_{S}$ be the shape’s orientation, and suppose $f^{*}:\mathcal{S}\rightarrow SO(3)$ is a solution to the following quotient regression problem:
+Let $\Omega_{S}$ be the shape’s orientation, and consider the following quotient regression problem:
 $\min_{f:\mathcal{S}\rightarrow SO(3)}\underset{R\sim U(SO(3))}{\mathbb{E}}\left[\min_{Q\in\hat{\mathcal{R}}}\|f(RS)-RQ\Omega_{S}\|_{F}^{2}\right],$ (4)
-Then for any $R\in SO(3)$, $f^{*}(RS)=RQ^{*}\Omega_{S}$ for some $Q^{*}\in\hat{\mathcal{R}}$.
+Then for any $R\in SO(3)$, there exists a solution $f^{*}:\mathcal{S}\rightarrow SO(3)$ of the form $f^{*}(RS)=RQ^{*}\Omega_{S}$ for some $Q^{*}\in\hat{\mathcal{R}}$.
 We prove this proposition in Appendix A.4.
@@
 We provide full architecture and training details in Appendix B.
-We also employ test-time augmentation to improve our flipper model’s predictions.
+We also employ test-time augmentation (TTA) to improve our flipper model’s predictions.
 Similarly to the case with the quotient orienter, we (1) randomly flip the inputs by $K$ random rotations $R_{k}\sim\hat{\mathcal{R}}=\mathcal{O}$, (2) obtain the flipper’s predictions for each shape, (3) return these predictions to the original input’s orientation, and (4) output the plurality prediction.
@@
 Shapenet ModelNet40
-Ours 89.2 % 77.7 %
+Ours (with TTA) 89.2 % 77.7 %
+Ours (without TTA) 85.3 % 72.3 %
 Upright-Net (Pang et al., 2022) 69.5 % 62.3 %
@@
 “Upright-Net-Oracle” upper-bounds the performance of a hypothetical orienter built on Upright-Net, whereas “Upright-Net-Random” provides a lower bound on such an orienter’s performance.
-We benchmark our pipeline’s performance (“Orient Anything”) against each baseline in Table 2, where we report the mean chamfer distance between each model’s predicted shapes and the corresponding ground truth shapes, which are in canonical orientation.
-Orient Anything achieves an 84% reduction in mean chamfer distance between predicted and ground truth shapes relative to the oracle baseline, which augments Upright-Net’s predicted up-axis with the shape’s ground truth front-axis.
+We benchmark our pipeline’s performance against each baseline in Table 2, where we report the mean chamfer distance between each model’s predicted shapes and the corresponding ground truth shapes, which are in canonical orientation.
+Our method achieves an 84% reduction in mean chamfer distance between predicted and ground truth shapes relative to the oracle baseline, which augments Upright-Net’s predicted up-axis with the shape’s ground truth front-axis.
 In practice, one does not have access to this front-axis at inference time, so this represents a loose upper bound on the potential performance of a shape orientation algorithm built upon Upright-Net’s predicted up-axes.
@@
 Upright-Net-Oracle 0.05481 $\pm$ 0.13016
-Orient Anything 0.00856 $\pm$ 0.0396
-Orient Anything-APS 0.00208 $\pm$ 0.01407
+Ours (TTA) 0.00856 $\pm$ 0.0396
+Ours (w/o TTA) 0.01107 $\pm$ 0.04342
+Ours (with APS) 0.00208 $\pm$ 0.01407
 Figure 8: Histogram of APS sizes with coverage probability $\alpha=0.3$. The median APS contains 2 shapes, and 90% of APSs contain at most 8 shapes.
@@
 By applying each flip to the first-stage orienter’s output, we obtain a shape orientation pipeline that outputs sets of candidate shapes $\hat{S}_{c}$ given a single arbitrarily-rotated input $RS$.
-We call this method “Orient Anything-APS.” To evaluate its performance, we compute the chamfer distance between each candidate shape and the ground truth shape, and take the minimum over these chamfer distances.
+To evaluate its performance, we compute the chamfer distance between each candidate shape and the ground truth shape, and take the minimum over these chamfer distances.
 As one may trivially reduce the minimum chamfer distance over an APS by outputting arbitrarily large sets, we report a histogram of APS sizes in Figure 8 to demonstrate that these sets are typically small.
-We report the performance of “Orient Anything-APS” in the final row of Table 2.
+We report our pipeline’s performance with adaptive prediction sets in the last row of Table 2.
 Allowing our pipeline to output an APS reduces the mean chamfer distance between the closest shape in this set to the ground truth by a factor of 4 relative to Orient Anything without this feature.
@@
 We anticipate that this factorization of geometric learning problems will be broadly applicable throughout 3D deep learning for tackling problems that are ill-posed due to the presence of symmetries.
-Impact statement.
-This paper presents a novel algorithm for 3D shape orientation. We do not believe that there are notable ethical considerations or societal impacts of our work meriting specific consideration.
-Acknowledgments.
+Acknowledgments
 We thank the authors of Pang et al.
 (2022) for their gracious assistance with their source code and Justin Solomon for valuable discussions.
-CS thanks Backflip AI for hosting him and generously supporting him with mentorship and compute throughout his internship.
+Chris thanks Backflip AI for hosting him and generously supporting him with mentorship and compute throughout his internship, and gratefully acknowledges the support of a 2024 Exponent fellowship.
+Impact statement
+This paper presents a novel algorithm for 3D shape orientation. We do not believe that there are notable ethical considerations or societal impacts of our work meriting specific consideration.
 Appendix A Additional results and proofs
@@
 We train the quotient orienter for 1919 epochs at a learning rate of $10^{-4}$.
-For test-time augmentation, we (1) randomly rotate the inputs $RS$ by $K$ random rotations $R_{k}\sim U(SO(3))$, $k=1,...,K$, (2) obtain the quotient orienter’s predictions $f_{\theta}(R_{k}RS)$ for each shape, (3) return these predictions to the original input’s orientation by computing $R_{k}^{\top}f_{\theta}(R_{k}RS)$, and (4) output the prediction $R_{k^{*}}^{\top}f_{\theta}(R_{k^{*}}RS)$ with the smallest average quotient distance to the remaining predictions.
+Test-time augmentation for the quotient orienter.
+Our use of TTA for the quotient orienter is motivated by the observation that the quotient orienter succeeds on most rotations and fails on small subsets of rotations.
+Given an arbitrarily-oriented input shape $RS$, we would like to mitigate the possibility that the shape is in an orientation for which the quotient orienter fails.
+To do so, we apply $K$ random rotations $R_{k}$ to the input shape to obtain randomly re-rotated shapes $R_{k}RS$.
+Because the quotient orienter only fails on small subsets of rotations, we expect it to succeed on most of the $R_{k}RS$ and recover the orientations of $R_{k}RS$ up to an octahedral symmetry.
+Because we are interested in the orientation of $RS$ rather than $R_{k}RS$, we then need to apply the inverse of $R_{k}$ to each predicted orientation $f_{\theta}(R_{k}RS)$ to obtain a set of $K$ predicted orientations $R_{k}^{\top}f_{\theta}(R_{k}RS)$.
+These will be correct orientations of $RS$ up to octahedral symmetries for each $k$ where the quotient orienter succeeds.
+Because we expect this to be the case for most $k$, we employ a voting scheme to select one of the candidate orientations $R_{k}^{\top}f_{\theta}(R_{k}RS)$.
+In this step, we compute each candidate’s average quotient $L_{2}$ loss with respect to every other candidate (i.e.
+the loss in Problem 3), and choose the candidate which minimizes this measure.
+Because we expect the quotient orienter to have succeeded for most of the $R_{k}RS$, this average quotient loss will be small for most candidates and large for the few outlier candidates on which the orienter failed.
+Choosing the candidate with the minimum average quotient loss with respect to the other candidates filters out these outliers and makes it likelier that we output the correct orientation of $RS$ up to an octahedral symmetry.
 B.2 Flipper
@@
 We draw rotations $Q\in U(\mathcal{O})$ during training, and simulate inaccuracies in our quotient orienter’s predictions by further rotating the training shapes about a randomly drawn axis by an angle uniformly drawn from $[0,10]$ degrees.
-We also employ test-time augmentation to improve our flipper model’s predictions.
-Similarly to the case with the quotient orienter, we (1) randomly flip the inputs by $K$ random rotations $R_{k}\sim\hat{\mathcal{R}}=\mathcal{O}$, (2) obtain the flipper’s predictions for each shape, (3) return these predictions to the original input’s orientation, and (4) output the plurality prediction.
+Test-time augmentation for the flipper.
+The TTA procedure for the flipper is similar to the quotient orienter’s TTA procedure.
+We are given an input shape $FS$, which is correctly oriented up to an octahedral symmetry (a “flip”) $F$ if the quotient orienter succeeded.
+We apply $K$ random flips $F_{k}$ to the input shape to obtain randomly re-rotated shapes $F_{k}FS$ and apply the flipper to each of these shapes.
+Similarly to the quotient orienter, if the flipper succeeds on some $F_{k}FS$, it predicts $F_{k}F$ rather than the flip $F$ we are actually interested in.
+We consequently left-multiply each prediction $g(F_{k}FS)$ by $F_{k}^{\top}$, which maps each successful prediction $g(F_{k}FS)=F_{k}F$ to the true flip $F$.
+Because we expect the flipper to have succeeded on most inputs $F_{k}FS$ and failed on a minority of inputs, we again use a voting scheme to pick out the plurality prediction.
+In this case, we simply return the most common flip among the set of $F_{k}^{\top}g(F_{k}FS)$.
 B.3 Adaptive prediction sets
@@
 Conversely, the sets will be large when the flipper is uncertain and assigns similar mass to most classes.
+B.4 Motivation for test-time augmentation.
+In practice, our trained orienter exhibits discontinuities at certain rotations.
+To demonstrate this, we perform an experiment on a bench shape, which is symmetric under a 180 degree rotation about the y-axis.
+We rotate the bench around this axis in increments of 1 degree and pass the resulting shape into our orienter for each rotation.
+We track two metrics throughout this process:
+  • The quotient loss of our orienter’s prediction (i.e.
+the loss function from Problem 3, where we quotient the $L_{2}$ loss by the octahedral group).
+This measures the accuracy of our orienter’s predictions up to an octahedral symmetry.
+  • The chamfer distance between the shapes obtained by applying the orienter’s output to the rotated bench at subsequent angles of rotation. This detects discontinuities in the orienter’s output.
+We plot both of these metrics across the rotation angle in Figure 9.
+Our orienter exhibits several sharp discontinuities, manifested as spikes in the chamfer distance plot.
+These discontinuities are associated with spikes in the quotient loss, which are occasionally large.
+However, these spikes in the quotient loss are highly localized, and our orienter performs well for the vast majority of rotations.
+These localized spikes motivate our use of TTA to improve our pipeline’s performance.
+We provide implementation details for TTA in Appendices B.1 and B.2.
+Figure 9: Our orienter exhibits several sharp discontinuities when rotating an input shape about a fixed axis (red curve).
+These discontinuities are associated with localized spikes in the quotient loss (blue curve), which motivate our use of TTA to improve our pipeline’s performance.
 Appendix C Additional figures